### v1

# Instruction-tuning Aligns LLMs to the Human Brain

Abstract
Instruction-tuning is a widely-adopted method of finetuning that enables large language models (LLMs) to generate output that more closely resembles human responses to natural language queries, in many cases leading to human-level performance on diverse testbeds. However, it remains unclear whether instruction-tuning truly makes LLMs more similar to how humans process language. We investigate the effect of instruction-tuning on LLM-human similarity in two ways: (1) brain alignment, the similarity of LLM internal representations to neural activity in the human language system, and (2) behavioral alignment, the similarity of LLM and human behavior on a reading task. We assess 25 vanilla and instruction-tuned LLMs across three datasets involving humans reading naturalistic stories and sentences. We discover that instruction-tuning generally enhances brain alignment by an average of $6\%$, but does not have a similar effect on behavioral alignment. To identify the factors underlying LLM-brain alignment, we compute correlations between the brain alignment of LLMs and various model properties, such as model size, various problem-solving abilities, and performance on tasks requiring world knowledge spanning various domains. Notably, we find a strong positive correlation between brain alignment and model size (r = 0.95), as well as performance on tasks requiring world knowledge (r = 0.81). Our results demonstrate that instruction-tuning LLMs improves both world knowledge representations and brain alignment, suggesting that mechanisms that encode world knowledge in LLMs also improve representational alignment to the human brain.
1 Introduction
Instruction-tuning is a widely adopted method for finetuning large language models (LLMs) on datasets containing task-specific instructions. This approach enhances their ability to generalize effectively to previously unseen tasks by learning to follow provided instructions (Wang et al., 2022c). Instruction-tuning often costs only a small fraction of compute relative to pretraining (Chung et al., 2022), yet propels pretrained LLMs to incredible performance leaps on reasoning and problem-solving benchmarks. This transformation has enabled LLMs to approach human performance on many tasks, despite using only few (or zero) training examples, as well as to tackle open-world reasoning tasks previously only achievable by humans (Zhang et al., 2023).
In addition to teaching LLMs to understand and follow human instructions, instruction-tuning also improves the ability of LLMs to mimic the ground-truth outputs (often human-written) of the training data. This property allows them to produce more controllable and predictable output that is deemed (1) more desirable by human evaluators on various metrics (Zhang et al., 2023; Chung et al., 2022; Wang et al., 2022b), (2) more aligned to human values (Chia et al., 2023), and (3) more stylistically similar to human outputs (Dasgupta et al., 2022; Safdari et al., 2023).
Consequently, instruction-tuning yields LLMs more similar to humans in both capability and output rememblance. From a neuroscience perspective, this begs the question: Does instruction-tuning make LLMs more similar to the human language system? Previous work has shown that models with higher task performance are more aligned to the human language system (Schrimpf et al., 2021; Goldstein et al., 2022; Caucheteux & King, 2022), and, on some datasets, even hit the estimated noise ceiling. [^1] In fMRI recordings, an upper limit of representational similarity can be computed by sampling from the same participant twice, deducing a threshold defined by the noise level of the data gathering process. However, there has been no similar study on how instruction-tuning, the method that has enabled powerful LLMs such as ChatGPT, affects alignment to the human language system.
In this work, we explore the impact of instruction-tuning on the alignment between LLMs and the human language system, considering two aspects: (1) brain alignment, which assesses how closely LLMs’ internal representations match neural activity patterns in the human language system, and (2) behavioral alignment, which evaluates the similarity between LLM behavior and human behavior. To conduct this study, both LLMs and human participants are presented with the same language stimuli comprised of naturalistic stories and sentences. For LLMs, we analyze their internal representations and per-word perplexity, while for humans, we use previously collected brain activity data from functional magnetic resonance imaging (fMRI) experiments and per-word reading times.
To measure brain alignment, we use the Brain-Score (Schrimpf et al., 2018) linear predictivity metric, assessing how well LLM representations predict human brain activity in response to the same language stimuli (Jain & Huth, 2018; Toneva & Wehbe, 2019; Schrimpf et al., 2021; Oota et al., 2023), using data from three neural datasets: Pereira et al. (2018), Blank et al. (2014), and Wehbe et al. (2014). To evaluate behavioral alignment, we use a benchmark in Brain-Score which calculates the Pearson correlation between LLM per-word perplexity and human per-word reading times from the Futrell et al. (2018) dataset. Perplexity for LLMs and reading times for humans offer insights into comprehension difficulty (Ehrlich & Rayner, 1981; Hale, 2001; Smith & Levy, 2013), allowing us to examine whether LLMs and humans share similarities in terms of which words and sentences they find challenging or surprising. Because models vary in their brain and behavioral alignment across different architectures and training objectives (Schrimpf et al., 2021), we estimate the effect of instruction-tuning across 17 instruction-tuned LLMs and 8 vanilla LLMs, and report a significant increase in brain alignment by instruction-tuned models compared to vanilla ones.
To investigate why instruction-tuning increases alignment to human brain activity, we then estimate the contribution of various LLM properties towards brain alignment. Specifically, we compute Pearson correlations between an LLM’s brain alignment and its properties, including next-word prediction (NWP) ability, model size, a range of problem-solving abilities, and world knowledge spanning different domains. The evaluation of the latter two properties is based on the Big-Bench Hard benchmark (BBH) (Suzgun et al., 2022) and the Massive Multi-task Language Understanding benchmark (MMLU) (Hendrycks et al., 2021), respectively.
We report three major findings:
  1. 1.
    Instruction-tuning generally improves the alignment of LLM representations to brain activity, increasing brain alignment by 6.26.2% on average for the LLMs and neural datasets we tested (Figure 1).
  2. 2.
    Investigating the factors underlying LLM-brain alignment, we find that world knowledge and model size are strongly correlated with brain alignment (r = $0.81$ and $0.95$ for instruction-tuned models, respectively; Figure 2).
  3. 3.
    Surprisingly, our results generally indicate that instruction-tuning LLMs does not enhance behavioral alignment with human reading times. Furthermore, behavioral alignment on this dataset demonstrates poor correlations with all other measures we investigate, including task performance and model size (Figure 3).
Figure 1: Instruction-tuning aligns LLM representations to human brain activity. (A) We evaluate brain alignment as the similarity of an LLM’s internal representations to human brain activity, using a linear predictivity metric in Brain-Score. We evaluate 25 vanilla and instruction-tuned LLMs with sizes between 77M and 33B parameters. We use 3 neural datasets of humans reading naturalistic stories and sentences: Pereira2018, Blank2014, and Wehbe2014, and compute the average across the three datasets. (B) Instruction-tuning improves average brain alignment by 6.2% on average. Each point above the identity line represents an instruction-tuned LLM that has greater brain alignment than its vanilla counterpart. Error bars (here and elsewhere) represent median absolute deviation over human participants. (C) Instruction-tuning generally improves brain alignment on all three neural datasets. (D) We instruction-tune LLaMA-7B using the Alpaca dataset (“Instruction” model). We also train an ablation model with the same process and training data, but remove the instruction portion from each training sample (“No Instruction” model). This experiment shows that brain alignment improvements are due to both (1) training data (present in both models) and (2) the process of training LLMs to understand and follow instructions (present only in original model).
2 Background & Related Work
Effect of Instruction-tuning on LLMs.   Instruction-tuning is an effective method for enhancing LLM capability and controllability. It trains LLMs using pairs of human instructions and desired outputs. The benefits of instruction-tuning can be categorized into three key aspects (Zhang et al., 2023): (1) it bridges the disparity between the pretraining objective of LLMs (next-word prediction) and the goal of accurately following human instructions, (2) it achieves greater control and predictability of model behavior compared to standard LLMs, allowing researchers to make them more similar to humans in both capability and output similarity (Chia et al., 2023; Dasgupta et al., 2022; Safdari et al., 2023), and (3) it often costs only a small fraction of compute relative to pretraining, enabling LLMs to swiftly adapt to target domains (Chung et al., 2022). We contribute to this research area from a neuroscience perspective, by studying whether instruction-tuning makes LLMs more aligned to the human language system in terms of brain and behavioral alignment.
Effect of Finetuning on Brain alignment.   Prior works have studied how finetuning affects LMs’ alignment to human brain activity. These include finetuning on a wide range of downstream NLP tasks (Oota et al., 2022), finetuning to summarize narratives (Aw & Toneva, 2023), and finetuning to directly predict brain activity recordings (Schwartz et al., 2019). These studies aim to use brain alignment to study how finetuning affects LMs and their representations. Our work builds on this line of research by demonstrating that instruction-tuning aligns LLM representations to human brain activity. We also investigate why instruction-tuned LLMs align to brain activity by testing the correlation of brain alignment with various world knowledge domains and problem-solving abilities.
LM properties linked to Brain alignment.   There is a growing body of work disentangling the contribution of various LM properties towards brain alignment. These include studying how brain alignment is driven by next-word prediction ability (Schrimpf et al., 2021; Caucheteux & King, 2022), multi-word semantics (Merlin & Toneva, 2022), performance on various NLP tasks (Oota et al., 2022), and model size (Antonello et al., 2023). To disentangle the contribution of various LM properties toward brain alignment, we use instruction-tuned LLMs. They are especially useful as they have been trained to respond to a standard question-answer format, allowing us to evaluate LLMs on a wide array of tasks and in a more fine-grained manner. We expand this area of research by identifying that world knowledge is a key property underlying LLM-brain alignment.
3 Language Models
We evaluate the brain alignment of 25 large language models (LLMs) from two model families: T5 (Raffel et al., 2020) and LLaMa (Touvron et al., 2023). T5 models are encoder-decoder LLMs pre-trained on the Colossal Common Crawl Corpus (C4), a corpus of 356 billion tokens, using a masked infilling objective, and then further finetuned on multi-task mixture of unsupervised and supervised tasks converted into a text-to-text format. In our study, we use all five T5 models with sizes between 77M to 11B parameters. LLaMA models (Touvron et al., 2023) are decoder-only LLMs trained on 1.6 trillion tokens from a mixture of corpora including C4, English CommonCrawl, Wikipedia, Github, and more. For LLaMA, we use the 7B, 13B, and 33B parameter versions in our study.
For the instruction-tuned variants of T5 models, we utilize a variety of models finetuned on the FLAN suite (1515M examples for 1,836 different tasks accompanied by instructions, Chung et al., 2022), Alpaca (5252K instruction-following examples generated through methods inspired by Self-Instruct, Wang et al. (2022a), Taori et al., 2023), and GPT4ALL (437437K instruction-following examples generated with GPT-3.5-turbo, Anand et al., 2023) datasets. As for the LLaMa model family, we employ Vicuna’s 77B, 1313B, and 3333B models (Chiang et al., 2023), which are finetuned on user-shared conversations. Additionally, we incorporate the StableVicuna-13B model, which further refines the Vicuna-13B model using reinforcement learning from human feedback (RLHF) (Ouyang et al., 2022) on a range of conversational and instructional datasets. We also use the 77B version of Alpaca (Taori et al., 2023). Additional details about these LLMs can be found in Appendix A.
Table 1: Example of instruction-tuning training data format: (Instruction, Input, Output) from the Alpaca dataset (Taori et al., 2023). The input field is optional for certain types of instructions.
Instruction Input Output
“Write a short paragraph “The importance of “The use of renewable energy is growing
about the given topic.” using renewable energy.” rapidly in relevance and …”
4 Brain Alignment
Brain alignment refers to the method of evaluating the similarity between LLM representations and human brain activity (Figure 1). This relies on fMRI recordings of human subjects while they read language stimuli on potentially any topic (here: Pereira et al., 2018; Blank et al., 2014; Wehbe et al., 2014). The same language stimuli from prior brain recordings are provided as input to LLMs, whose intermediate layer activations are extracted as their representations of the language stimuli. We follow a general approach previously used in several works (Schrimpf et al., 2018; 2021; Jain & Huth, 2018; Toneva & Wehbe, 2019; Oota et al., 2023; Aw & Toneva, 2023). Specifically, we use the linear predictivity metric implemented in Brain-Score (Schrimpf et al., 2020, Figure 1), first training a linear function to predict fMRI voxels associated with the human language system using LLM representations as input features. We then apply this linear function to held-out brain activity data from the original corpus of recordings, and evaluate the brain alignment of the LLM as the Pearson correlation between the predicted and actual brain activity data. For each LLM, we evaluate its brain alignment for every layer (e.g., LLaMA-7B has 32 layers), and use the highest value as the LLM’s brain alignment value, following Schrimpf et al. (2018).
Datasets   We use three fMRI datasets to measure the brain alignment of LLMs. Each dataset involves a different set of human participants and uses a different set of language stimuli.
Pereira2018 (experiments 2 and 3 from Pereira et al., 2018): In experiment 2, nine participants read 384 sentences organized into 96 text passages. In experiment 3, six participants read 243 sentences in 72 text passages. Each sentence was displayed for four seconds on a screen.
Blank2014 (Blank et al., 2014): The data consists of fMRI recordings of 5 human participants listening to naturalistic stories from the Natural Stories Corpus (Futrell et al., 2018).
Wehbe2014 (Wehbe et al., 2014): The data includes fMRI recordings of 8 human participants reading chapter 9 of the book Harry Potter and the Sorceror’s Stone (Rowling et al., 1998). Participants read the chapter at a fixed interval of one word every 0.5 seconds.
4.1 Instruction-tuning aligns LLM representations to human brain activity
First, we study the effect of instruction-tuning on LLM brain alignment. We compute each LLM’s average brain alignment as the mean of its brain alignment on the 3 neural datasets. We find that instruction-tuning improves alignment by an average of 6.2% across all tested LLMs (Figure 1B). This holds across all three neural datasets, with average improvements of +6.9% on Pereira2018, +8.0% improvement on Blank2014, and +3.8% on Wehbe2014 (Figure 1C). This provides a strong signal that instruction-tuning improves brain alignment. Moreover, a smaller instruction-tuned model can attain higher brain alignment than a larger model from the same family that has not been instruction-tuned (e.g., Alpaca-7B v.s. LLaMa-13B, see detailed results in Appendix E).
To longitudinally study how instruction-tuning aligns LLM representations to brain activity, we separately instruction-tune a LLaMA-7B model on the Stanford Alpaca instruction dataset (Taori et al., 2023) for 3 epochs. By evaluating checkpoints regularly during training, we find that instruction-tuning progressively improves brain alignment (Figure 1D). We also perform an ablation study to disambiguate the effect on brain alignment of (1) the instruction-following ability provided by the instruction-tuning step and (2) added training data. We fine-tune LLaMA-7B with the same process and data, but remove the instruction portion from each training sample. We observe that brain alignment of this ablated model increases during fine-tuning but stays lower than its instruction-following counterpart (Figure 1D). This shows that brain alignment improvements are due to both factors.
Figure 2: World knowledge and model size are important factors underlying LLM-brain alignment. To identify factors underlying brain alignment, we test Pearson correlations between brain alignment and various LLM properties, such as model size, world knowledge in various domains (MMLU benchmark), and various types of problem-solving abilities (BBH benchmark). Insets display results on individual datasets, with stars reflecting statistical significance (n.s. = p >> 0.05, * = p << 0.05, ** = p << 0.005, etc.) (A) Brain alignment is significantly and strongly correlated with world knowledge as evaluated by the MMLU Overall score (r = 0.81), which reports the mean performance across all MMLU subjects. (B) Brain alignment is significantly and strongly correlated with performance on the world knowledge task category in BBH (r = 0.68). (C) Brain alignment is significantly and strongly correlated with model size (logarithm of number of parameters) (r = 0.95).
4.2 Factors underlying LLM-brain alignment
To identify factors underlying the representational similarity between LLMs and human brains, we compute the Pearson correlation between LLM brain alignment and various properties of LLMs: performance on a benchmark involving different reasoning abilities (BBH benchmark; Suzgun et al., 2022), performance on a benchmark requiring domain-specific world knowledge (MMLU; Hendrycks et al., 2021), language modeling ability, and model size.
World Knowledge and Reasoning abilities   MMLU is designed to measure the world knowledge of LLMs across many subject domains. It contains 57 tasks, categorized by the subject domain of world knowledge tested: STEM, Humanities, Social Sciences, and Others (a broad category ranging from finance to marketing to professional medicine). BBH contains 23 tasks, grouped into four categories based on the type of problem-solving ability tested: Algorithmic and Multi-Step Arithmetic Reasoning; Natural Language Understanding; Use of World Knowledge; and Multilingual Knowledge and Reasoning. For both benchmarks, we use category classifications from the original papers. We measure the performance of LLMs on BBH and MMLU using the instruct-eval repository [^2] https://github.com/declare-lab/instruct-eval with default settings (3-shots, 5-shots respectively) and preset prompts. We measure the Pearson correlation (and its p-value) between LLM brain alignment and performance on each category of the MMLU and BBH benchmarks. We obtain p-values after performing false discovery rate correction, ensuring our results are significant and not due to random chance.
We find that brain alignment is significantly and strongly correlated with world knowledge. On the MMLU benchmark, we observe a high correlation between brain alignment scores and the MMLU Overall score (r = 0.81), which reports the mean performance across all world knowledge subject domains on MMLU (Figure 2A). Similarly, brain alignment is also significantly and strongly correlated with the mean performance on tasks included in the world knowledge category of the BBH benchmark (r = 0.68; Figure 2B). Interestingly, we do not find strong correlations with other dimensions of the BBH benchmark (e.g., Algorithmic reasoning and Multilingual reasoning, see Table 2), though this could also be due to limitations of the tested models (most are primarily pretrained on English language data), as indicated by their low raw performance scores on some tasks. Overall, our results provide a strong signal that more accessible representations of world knowledge are a key factor in aligning LLM representations to human brain activity.
Language Modeling Ability   Prior works have shown correlations between brain alignment and next-word prediction (NWP) ability (Caucheteux & King, 2022; Schrimpf et al., 2021). We find similar results for correlation between brain alignment and NWP loss (r = -0.54, Appendix G). Interestingly, the strength of correlation between brain alignment and world knowledge performance (r = 0.81) is greater than that between brain alignment and NWP loss (r = -0.54). This result suggests that world knowledge understanding is a better predictor of brain alignment than NWP ability.
Model Size   Finally, we find that brain alignment is significantly and strongly correlated with model size (r = 0.95), as measured by the logarithm of the number of model parameters (Figure 2C). Schrimpf et al. (2021) observe such a pattern for language models, and we find the pattern holds for instruction-tuned models, and models trained at a larger scale than their study (7B+ parameters).
However, model size alone does not determine brain alignment. Our results show that smaller instruction-tuned LLMs can have greater brain alignment than larger vanilla models. For example, LLaMA-13B obtains brain alignment of 0.220, Vicuna-13B obtains 0.229, LLaMA-33B obtains 0.227, and Vicuna-33B obtains 0.232. Hence, Vicuna-13B has greater brain alignment than LLaMA-33B despite being less than 40% its size. We observe a similar trend in another four models: T5-base, Flan-T5-base, T5-large, Flan-T5-large. Also, prior works have shown that large random models achieve poor brain alignment (Schrimpf et al., 2021). These results demonstrate that there are LLM properties aside from model size that contribute significantly to brain alignment.
Datasets   We observe significant correlations between brain alignment and world knowledge and model size on Pereira2018 and Wehbe2014, but surprisingly not Blank2014. We hypothesize that it may be because Blank2014 has a small participant size (N = 5) that leads to greater noise in the results, compared to Pereira2018 and Wehbe2014.
Table 2: Brain alignment strongly correlates with world knowledge across all subject domains in MMLU, as well as the world knowledge category in BBH. At the same time, brain alignment is not significantly correlated with all other types of problem-solving abilities in BBH (e.g., algorithmic or multilingual reasoning). We obtain p-values after performing false discovery rate correction, ensuring our results are significant and not due to random chance.
Task category Brain Alignment corrected Number Average Model
Correlation (r) p-value of tasks Performance
MMLU – Overall Score 0.809 0.000329 57 0.36
MMLU – STEM 0.792 0.000343 18 0.28
MMLU – Humanities 0.791 0.000343 13 0.34
MMLU – Social Sciences 0.807 0.000329 12 0.41
MMLU – Others 0.809 0.000329 14 0.40
BBH – Overall score 0.384 0.177 23 0.28
BBH – Algorithmic reasoning 0.194 0.558 8 0.22
BBH – Language understanding 0.163 0.585 3 0.43
BBH – World knowledge 0.679 0.005 5 0.36
BBH – Multilingual reasoning -0.035 0.895 1 0.19
BBH – Others 0.478 0.083 6 0.27
5 Behavioral Alignment
Figure 3: Instruction-tuning LLMs generally does not improve behavioral alignment to human reading times. Furthermore, behavioral alignment correlates poorly with all other tested measures: world knowledge, model size, and next-word prediction (NWP) ability. (A) We present the same language stimuli to LLMs and human participants, using the Futrell2018 benchmark in Brain-Score, which contains naturalistic stories. We compute the behavioral alignment as the Pearson correlation between LLM perplexity for each word and human reading times for each word. (B) Instruction-tuning does not generally improve behavioral alignment. Furthermore, behavioral alignment is poorly and not significantly correlated with all other measures: (C) world knowledge on MMLU (r = 0.08, p = 0.76), (D) model size (r = 0.26, p = 0.31), (E) NWP loss for T5 models (r = -0.2, p = 0.54), and (F) NWP loss for LLaMA models (r = 0.68, p = 0.21).
In the previous section, we show that instruction-tuning aligns the internal representations of LLMs to human brain recordings (Section 4.1). In this section, we explore whether instruction-tuning also aligns LLM behavior to human behavior.
Following the approach previously proposed by Schrimpf et al. (2021) and implemented in the Brain-Score package (Schrimpf et al., 2020), we measure behavioral alignment by evaluating the similarity between LLM per-word perplexity and human per-word reading times, given the same language stimuli (Figure 3A). We use the self-paced reading times dataset from Futrell et al. (2018), consisting of the reading times of 179 human participants recorded while they were visually presented with 10 naturalistic stories. We provide language stimuli from this data as input to LLMs and measure their per-word perplexity. Finally, we evaluate behavioral alignment by computing the Pearson correlation between per-word LLM perplexity and per-word human reading times.
Using the same set of models as in the previous section, we compare the behavioral alignment of each instruction-tuned LLM against its vanilla counterpart. Our results generally indicate that instruction-tuning LLMs does not improve behavioral alignment to human reading times (Figure 3B). For half of the LLMs tested, it results in no change or reduced behavioral alignment. Then, we test the correlation between LLM behavioral alignment and model size, next-word prediction ability, various reasoning abilities (as measured by performance on the BBH benchmark), and world knowledge across various domains (as measured by performance on the MMLU benchmark). Contrary to our findings on the correlation between brain alignment and model size and world knowledge (Section 4.2), we do not find that these factors are correlated with the behavioral alignment of models: world knowledge in Figure 3C (r = 0.08, p = 0.76), model size in Figure 3D (r = 0.26, p = 0.31), next-word prediction loss for T5 models in Figure 3E (r = -0.2, p = 0.54), and next-word prediction loss for LLaMA models in Figure 3F (r = 0.68, p = 0.21). We discuss our interpretations of these results and possible explanations in Section 6.2.
6 Discussion
6.1 Implications for NLP: Building LLMs
Using brain alignment to interpret LLMs. Representations in the human brain support a wide range of downstream tasks that current models still struggle on. Prior works have used brain activity results to interpret neural networks (Dong & Toneva, 2023) and build more performant models (Dapello et al., 2020; 2022; Safarani et al., 2021). Instruction-tuning has emerged as a breakthrough technique to improve LLM abilities, quality of outputs, and allow LLMs to adapt to new tasks with minimal task-specific training. However, the manner in which instruction-tuning alters the internal representations of LLMs to achieve these improvements remains an open question. Brain activity data provides a neuroscientific angle to investigate this question. Our results show that instruction-tuning LLMs improves both world knowledge representations and human brain alignment. They suggest that instruction-tuning improves the ability of an LLM to retrieve world knowledge stored in its model weights and that the mechanisms that encode world knowledge in LLMs also improve representational alignment to the human brain. Overall, our results highlight the importance of integrating world knowledge in the development of future LLMs.
6.2 Implications for Neuroscience: Studying LLM-Human Alignment
Instruction-tuned LLMs are useful for studying LLM properties underlying brain and behavioral alignment.   To identify why LLM and human brains exhibit representational similarities, prior work has mostly focused on high-level properties such as model size (Antonello et al., 2023), and external behaviors such as predicting missing words Schrimpf et al. (2021); Caucheteux & King (2022). However, a key to understanding these similarities is to identify lower-level or internal properties of LLMs that underlie brain alignment. This includes the amount of knowledge LLMs contain, e.g., factual (AlKhamissi et al., 2022) and commonsense (Sap et al., 2020; Bosselut et al., 2019). Our work shows that we can harness instruction-tuned LLMs for this purpose as they have been trained to respond to a general question format, allowing us to evaluate LLMs in a more fine-grained manner. This allows the study of both internal (e.g., knowledge) and external (e.g., behavior) properties of LLMs, and how they correlate with brain and behavioral alignment.
Examining more dimensions of behavior.   To evaluate behavioral alignment, our work and many prior works compare LM and human next-word surprisal on reading tasks (Wilcox et al., 2020; Schrimpf et al., 2021; Eghbal A. Hosseini et al., 2023). This evaluates only a single dimension of LM and human behavior (per-word perplexity and reading times). On the models we test here, behavioral alignment is not significantly correlated with model size, world knowledge, or next-word prediction ability. While next-word prediction performance correlates with alignment to human reading times across a broad band of models (Schrimpf et al., 2021), this trend does not hold up in recent Transformer-based LMs (Oh & Schuler, 2023), having a surprising negative correlation with parameter count (Oh et al., 2022). Our results highlight the need to create more benchmarks to expand the dimensions of behavior examined for both LLMs and humans, in order to holistically evaluate LLM behavior, as well as LLM-human behavioral alignment.
Brain alignment datasets with humans performing diverse tasks.   Our work studies brain alignment using neural datasets that are limited to humans reading naturalistic stories and sentences in English. Unfortunately, there does not exist brain activity data for human participants attempting the BBH and MMLU benchmarks. This may explain why brain alignment is not significantly correlated with many categories of problem-solving on BBH, e.g., language understanding. In the future, we hope to study brain alignment with human participants performing more diverse sets of tasks, e.g., reading computer program code (Ivanova et al., 2020). This can identify more factors underlying LLM-brain alignment, and provide insights into how brain activity and the human language system may be shaped by various forms of problem-solving. Furthermore, for the neural datasets in our work, many of the larger models exceed the noise ceiling estimates (Appendix E), highlighting the need for more neural datasets (with better ways of computing noise ceiling estimates).
World knowledge shapes brain activity.   Our results show that world knowledge is a key factor in aligning LLM representations to human brain activity. LLMs with greater world knowledge across all tested subject domains produce representations that align more closely to human brain activity. Analogously, this suggests that world knowledge influences human brain activity, shaping the language comprehension systems in the brain.
6.3 Limitations and Future Work
Limited model families.   A model’s performance on MMLU and BBH is affected not only by the quantity of knowledge it possesses, but also by its ability to follow each benchmark’s instruction format. Hence, we restricted our study to two model families (T5 and LLaMA). This attempts to ensure that the models are trained on a similar instruction format, so they have similar ability to follow the instruction format of MMLU and BBH. Thus, any differences in their MMLU or BBH performance would point to differences in the quantity of knowledge they possess. However, we hope to add a greater variety of LLMs in future work to strengthen the generalizability of our claims.
Computational cost.   We investigate 25 LLMs, with the largest models having 33B parameters, on many datasets: brain alignment (Pereira2018, Blank2014, Wehbe2014), behavioral alignment (Futrell2018), next-word prediction (WikiText-2) and other evaluations (BBH, MMLU). Unfortunately, this makes it computationally expensive to replicate our experiments. In the hopes of addressing this concern, we provide our full results in Appendices E, F, G, H, and I.
7 Conclusion
We investigate whether instruction-tuning improves the alignment of LLMs to the human language system. We evaluate 25 LLMs with parameter sizes ranging from 77 million to 33 billion, across three neural datasets of humans reading naturalistic stories and sentences. We find that instruction-tuning generally improves the alignment of LLM representations to brain activity. Delving into the factors underlying LLM-brain alignment, we discover that world knowledge and model size are key determinants of brain alignment. This suggests that world knowledge helps shape representations in the human language system, and highlights the significance of integrating world knowledge in the development of future LLMs.
Reproducibility Statement
All models and code repositories utilized in this study are open-source and their corresponding links are provided in Appendix B and D for full transparency and reproducibility. For the calculation of brain and behavioral alignment, we relied on the Brain-Score repository (www.github.com/brain-score/language), which serves as a publicly accessible resource for conducting these assessments. We encourage researchers interested in replicating our findings to refer to the provided links and consult the Brain-Score repository for further details on datasets and the evaluation process. For measuring the performance of LLMs on BBH and MMLU benchmarks, we use the broadly-used instruct-eval repository (https://github.com/declare-lab/instruct-eval) with default settings.
Appendix A Language Models: Parameter count and Number of Layers
Table 3: Parameter count and number of layers for all 25 vanilla and instruction-tuned LLMs. The upper part contains encoder-decoder models of the T5 family, the lower parts decoder-only models of the LLaMA family. For the parameter count, “M” refers to million and “B” refers to billion. The number of layers for T5 models is a sum of the number of encoder and decoder layers.
Model Parameter Count Number of Layers
t5-small 77 M 16
flan-t5-small 77 M 16
t5-base 250 M 24
flan-t5-base 250 M 24
flan-alpaca-base 250 M 24
t5-large 800 M 48
flan-t5-large 800 M 48
flan-alpaca-large 800 M 48
t5-xl 3 B 48
flan-t5-xl 3 B 48
flan-alpaca-xl 3 B 48
flan-gpt4all-xl 3 B 48
flan-sharegpt-xl 3 B 48
flan-alpaca-gpt4-xl 3 B 48
t5-xxl 11 B 48
flan-t5-xxl 11 B 48
flan-alpaca-xxl 11 B 48
llama-7b 7 B 32
alpaca-7b 7 B 32
vicuna-7b 7 B 32
llama-13b 13 B 40
vicuna-13b 13 B 40
stable-vicuna-13b 13 B 40
llama-33b 33 B 60
vicuna-33b 33 B 60
Appendix B Language Models: Links to models weights
Table 4: Link to model weights for all 25 vanilla and instruction-tuned LLMs. The upper part contains encoder-decoder models of the T5 family, the lower parts decoder-only models of the LLaMA family. We provide these links for reproducibility purposes.
Model Link to model weights
t5-small www.huggingface.co/google/t5-v1_1-small
flan-t5-small www.huggingface.co/google/flan-t5-small
t5-base www.huggingface.co/google/t5-v1_1-base
flan-t5-base www.huggingface.co/google/flan-t5-base
flan-alpaca-base www.huggingface.co/declare-lab/flan-alpaca-base
t5-large www.huggingface.co/google/t5-v1_1-large
flan-t5-large www.huggingface.co/google/flan-t5-large
flan-alpaca-large www.huggingface.co/declare-lab/flan-alpaca-large
t5-xl www.huggingface.co/google/t5-v1_1-xl
flan-t5-xl www.huggingface.co/google/flan-t5-xl
flan-alpaca-xl www.huggingface.co/declare-lab/flan-alpaca-xl
flan-gpt4all-xl www.huggingface.co/declare-lab/flan-gpt4all-xl
flan-sharegpt-xl www.huggingface.co/declare-lab/flan-sharegpt-xl
flan-alpaca-gpt4-xl www.huggingface.co/declare-lab/flan-alpaca-gpt4-xl
t5-xxl www.huggingface.co/google/t5-v1_1-xxl
flan-t5-xxl www.huggingface.co/google/flan-t5-xxl
flan-alpaca-xxl www.huggingface.co/declare-lab/flan-alpaca-xxl
llama-7b www.github.com/facebookresearch/llama
alpaca-7b www.github.com/tatsu-lab/stanford_alpaca
vicuna-7b www.huggingface.co/lmsys/vicuna-7b-v1.3
llama-13b www.github.com/facebookresearch/llama
vicuna-13b www.huggingface.co/lmsys/vicuna-13b-v1.3
stable-vicuna-13b www.huggingface.co/CarperAI/stable-vicuna-13b-delta
llama-33b www.github.com/facebookresearch/llama
vicuna-33b www.huggingface.co/lmsys/vicuna-33b-v1.3
Appendix C MMLU and BBH benchmarks
MMLU is designed to measure knowledge from many domains Hendrycks et al. (2021). It contains 57 tasks, categorized by the subject domain of world knowledge tested: STEM, Humanities, Social Sciences, and Others. The STEM category includes questions on computer science, physics, mathematics, etc. The Humanities category includes questions on philosophy, law, history, etc. The Social Sciences category includes questions on politics, sociology, economics, geography, etc. The Others category includes questions on business topics such as finance, accounting, as well as general knowledge of global facts.
BBH is designed to evaluate various problem-solving and reasoning abilities of LLMs (Suzgun et al., 2022). BBH contains 23 tasks, categorized by the type of problem-solving ability tested: (1) Algorithmic and Multi-Step Arithmetic Reasoning, (2) Natural Language Understanding, (3) Use of World Knowledge, and (4) Multilingual Knowledge and Reasoning. The world knowledge category of BBH contains tasks that test for factual and general knowledge. Tasks requiring factual knowledge include: “Sports Understanding” and “Movie Recommendation”. Tasks requiring general knowledge include: “Causal Judgement”, which tests knowledge about causal-reasoning suppositions, and “Ruin Names”, which requires knowledge about human perception and usage of humor in the English language.
For both benchmarks, we adopt the same category classification as used in their original papers. We measure the performance of LLMs on BBH and MMLU using the instruct-eval repository [^3] https://github.com/declare-lab/instruct-eval with default settings (3-shots, 5-shots respectively) and preset prompts.
Appendix D Code Repositories
We use the Brain-Score repository to evaluate brain alignment for the Pereira2018 and Blank2014 datasets, as well as behavioral alignment for the Futrell2018 dataset. Link: www.github.com/brain-score/language.
We use an open-source repository to evaluate brain alignment for the Wehbe2014 dataset. Link: www.github.com/awwkl/brain_language_summarization, which builds on www.github.com/mtoneva/brain_language_nlp.
We use Instruct-Eval repository to evaluate MMLU and BBH scores. Link: www.github.com/declare-lab/instruct-eval.
We use Stanford Alpaca repository for instruction-tuning. Link: www.github.com/tatsu-lab/stanford_alpaca).
Appendix E Results for Brain alignment
Table 5: Brain alignment results for all 25 vanilla and instruction-tuned LLMs. We provide these results for reproducibility purposes.
Pereira2018 Blank2014 Wehbe2014 Average
t5-small 0.166 0.168 0.071 0.135
flan-t5-small 0.202 0.178 0.079 0.153
t5-base 0.222 0.188 0.074 0.162
flan-t5-base 0.234 0.178 0.076 0.163
flan-alpaca-base 0.227 0.179 0.076 0.161
t5-large 0.270 0.082 0.071 0.141
flan-t5-large 0.311 0.104 0.080 0.165
flan-alpaca-large 0.322 0.126 0.082 0.177
t5-xl 0.285 0.192 0.072 0.183
flan-t5-xl 0.314 0.215 0.072 0.200
flan-alpaca-xl 0.312 0.209 0.075 0.199
flan-gpt4all-xl 0.300 0.206 0.078 0.195
flan-sharegpt-xl 0.323 0.211 0.070 0.201
flan-alpaca-gpt4-xl 0.302 0.205 0.073 0.193
t5-xxl 0.343 0.297 0.096 0.246
flan-t5-xxl 0.350 0.268 0.103 0.240
flan-alpaca-xxl 0.346 0.268 0.102 0.239
llama-7b 0.405 0.154 0.118 0.226
alpaca-7b 0.420 0.167 0.118 0.235
vicuna-7b 0.399 0.152 0.119 0.223
llama-13b 0.412 0.133 0.115 0.220
vicuna-13b 0.423 0.148 0.116 0.229
stable-vicuna-13b 0.415 0.144 0.115 0.225
llama-33b 0.426 0.145 0.109 0.227
vicuna-33b 0.436 0.156 0.105 0.232
Table 6: Noise ceiling estimates for all 3 neural datasets. For Pereira2018 and Blank2014, noise ceiling estimates are computed using the Brain-Score repository, with details provided in Schrimpf et al. (2021). For Wehbe2014, noise ceiling estimates are also computed using a similar procedure.
Pereira2018 Blank2014 Wehbe2014 Average
Noise ceiling 0.359 0.210 0.104 0.224
Appendix F Results for Next-word prediction, MMLU, BBH
Table 7: WikiText-2 NWP loss, MMLU Overall Score, and BBH Overall Score for all instruction-tuned LLMs. Results for vanilla LLMs are not shown as they are not adapted for the question formats in the MMLU and BBH benchmarks. We provide these results for reproducibility purposes.
WikiText-2 NWP Loss MMLU Overall Score BBH Overall Score
flan-t5-small 0.851 0.294 0.287
flan-t5-base 1.235 0.341 0.308
flan-alpaca-base 1.074 0.304 0.266
flan-t5-large 0.625 0.419 0.370
flan-alpaca-large 0.648 0.397 0.276
flan-t5-xl 0.650 0.493 0.402
flan-alpaca-xl 0.604 0.466 0.270
flan-gpt4all-xl 0.625 0.337 0.212
flan-sharegpt-xl 0.664 0.446 0.363
flan-alpaca-gpt4-xl 0.593 0.456 0.348
flan-t5-xxl 0.638 0.545 0.443
flan-alpaca-xxl 0.607 0.508 0.229
alpaca-7b 4.201 0.404 0.328
vicuna-7b 4.387 0.472 0.331
vicuna-13b 4.130 0.521 0.387
stable-vicuna-13b 4.623 0.495 0.380
vicuna-33b 3.940 0.590 0.426
Notes on comparing next-word prediction (NWP) loss across model families.   The T5 and LLaMA models belong to separate model families. We wish to caution that comparing next-word prediction loss across different model families may not be meaningful. This is due to several reasons related to architectural differences, training methodologies, and objectives. (1) Architecture: T5 models have an encoder-decoder architecture while LLaMA models have a decoder-only architecture. (2) Training Objectives: The T5 models were trained on supervised and unsupervised tasks, while the LLaMA models were trained only on unsupervised text 3. (3) Loss computation: The loss functions for both model families are computed differently, making it inappropriate to directly compare their loss values. (4) Evaluation Metrics: Next-word prediction loss is just one metric, and it may not capture the overall language understanding capabilities of a model. Hence, we additionally evaluate these LLMs’ alignment to human brain activity, as well as their performance on problem-solving abilities (BBH) and tasks requiring world knowledge (MMLU). In summary, while NWP loss is a valuable metric for evaluating language models within the same family or architecture, comparing across different model families may not be meaningful.
Appendix G Results for Correlations of Brain Alignment with LLM properties
Figure 4: Correlation between brain alignment and various LLM properties: (A) MMLU benchmark global score, (B) BBH benchmark score with only world knowledge tasks, (C) number of parameters of the model, and (D) Next word prediction (NWP) performance.
Figure 5: Correlation between brain alignment and various LLM properties for all 25 LLMs: (A) number of parameters of the model, and (B) Next word prediction (NWP) performance.
Appendix H Results for Behavioral alignment
Table 8: Behavioral alignment results for all 25 vanilla and instruction-tuned LLMs. We provide these results for reproducibility purposes.
Futrell2018
t5-small 0.229
flan-t5-small 0.054
t5-base 0.333
flan-t5-base 0.152
flan-alpaca-base 0.290
t5-large 0.303
flan-t5-large 0.145
flan-alpaca-large 0.291
t5-xl 0.225
flan-t5-xl 0.113
flan-alpaca-xl 0.181
flan-gpt4all-xl 0.251
flan-sharegpt-xl 0.285
flan-alpaca-gpt4-xl 0.250
t5-xxl 0.260
flan-t5-xxl 0.274
flan-alpaca-xxl 0.267
llama-7b 0.204
alpaca-7b 0.206
vicuna-7b 0.205
llama-13b 0.184
vicuna-13b 0.184
stable-vicuna-13b 0.196
llama-33b 0.158
vicuna-33b 0.164
Table 9: Noise ceiling estimates for the Futrell2018 reading-times dataset. Noise ceiling estimates are computed using the Brain-Score repository, with details provided in Schrimpf et al. (2021).
Futrell2018
Noise ceiling 0.858
Appendix I Results for Instruction-tuning LLaMA-7B on Alpaca dataset
Figure 6: Improvements in brain alignment from instruction-tuning are due to both additional training data, as well as training to understand and follow instructions.
Instruction model
We instruction-tune LLaMA-7B on the Stanford Alpaca dataset (Taori et al., 2023) using the default training process, following the code in www.github.com/tatsu-lab/stanford_alpaca. In particular, the model is instruction-tuned using 5252K instruction-following examples generated through methods inspired by Self-Instruct (Wang et al., 2022a)). This model is labeled “Instruction” in Figure 6.
No instruction model (Ablation)
We also train an ablation model with the same process and training data as the default instruction-tuning, but remove the instruction portion from each training sample. This ablation model is labeled “No instruction” in Figure 6. This ablation experiment disentangles: (1) training data (present in both Instruction and No instruction), from (2) training LLMs to understand and follow instructions (present only in Instruction).
We use all provided training samples from the Alpaca dataset, thus ensuring that the models are trained on the same amount of data. We observe that brain alignment of the “No Instruction” ablation model increases during fine-tuning but stays lower than its “Instruction” counterpart.

### v2

@@
 Abstract
-Instruction-tuning is a widely-adopted method of finetuning that enables large language models (LLMs) to generate output that more closely resembles human responses to natural language queries, in many cases leading to human-level performance on diverse testbeds.
-However, it remains unclear whether instruction-tuning truly makes LLMs more similar to how humans process language.
-We investigate the effect of instruction-tuning on LLM-human similarity in two ways: (1) brain alignment, the similarity of LLM internal representations to neural activity in the human language system, and (2) behavioral alignment, the similarity of LLM and human behavior on a reading task.
-We assess 25 vanilla and instruction-tuned LLMs across three datasets involving humans reading naturalistic stories and sentences.
-We discover that instruction-tuning generally enhances brain alignment by an average of $6\%$, but does not have a similar effect on behavioral alignment.
-To identify the factors underlying LLM-brain alignment, we compute correlations between the brain alignment of LLMs and various model properties, such as model size, various problem-solving abilities, and performance on tasks requiring world knowledge spanning various domains.
+Instruction-tuning is a widely adopted finetuning method that enables large language models (LLMs) to generate output that more closely resembles human responses.
+However, no studies have shown that instruction-tuning actually teaches LLMs to process language in a similar manner as humans.
+We investigate the effect of instruction-tuning on aligning LLM and human language processing mechanisms in two ways: (1) brain alignment, the similarity of LLM internal representations to neural activity in the human language system, and (2) behavioral alignment, the similarity of LLM and human behavior on a reading task.
+We assess 25 vanilla and instruction-tuned LLMs on three datasets involving humans reading naturalistic stories and sentences, and find that instruction-tuning generally enhances brain alignment ($\sim$6%), but has no similar effect on behavioral alignment.
+To identify factors underlying this improvement in brain alignment, we compute correlations between brain alignment and various LLM properties, such as model size, problem-solving, and world knowledge understanding.
 Notably, we find a strong positive correlation between brain alignment and model size (r = 0.95), as well as performance on tasks requiring world knowledge (r = 0.81).
-Our results demonstrate that instruction-tuning LLMs improves both world knowledge representations and brain alignment, suggesting that mechanisms that encode world knowledge in LLMs also improve representational alignment to the human brain.
+Our results demonstrate that instruction-tuning LLMs improves both world knowledge representations and brain alignment, suggesting that the mechanisms that encode world knowledge in LLMs also improve representational alignment to the human brain.
 1 Introduction
-Instruction-tuning is a widely adopted method for finetuning large language models (LLMs) on datasets containing task-specific instructions.
-This approach enhances their ability to generalize effectively to previously unseen tasks by learning to follow provided instructions (Wang et al., 2022c).
-Instruction-tuning often costs only a small fraction of compute relative to pretraining (Chung et al., 2022), yet propels pretrained LLMs to incredible performance leaps on reasoning and problem-solving benchmarks.
-This transformation has enabled LLMs to approach human performance on many tasks, despite using only few (or zero) training examples, as well as to tackle open-world reasoning tasks previously only achievable by humans (Zhang et al., 2023).
-In addition to teaching LLMs to understand and follow human instructions, instruction-tuning also improves the ability of LLMs to mimic the ground-truth outputs (often human-written) of the training data.
-This property allows them to produce more controllable and predictable output that is deemed (1) more desirable by human evaluators on various metrics (Zhang et al., 2023; Chung et al., 2022; Wang et al., 2022b), (2) more aligned to human values (Chia et al., 2023), and (3) more stylistically similar to human outputs (Dasgupta et al., 2022; Safdari et al., 2023).
-Consequently, instruction-tuning yields LLMs more similar to humans in both capability and output rememblance.
-From a neuroscience perspective, this begs the question: Does instruction-tuning make LLMs more similar to the human language system?
-Previous work has shown that models with higher task performance are more aligned to the human language system (Schrimpf et al., 2021; Goldstein et al., 2022; Caucheteux & King, 2022), and, on some datasets, even hit the estimated noise ceiling.
-[^1] In fMRI recordings, an upper limit of representational similarity can be computed by sampling from the same participant twice, deducing a threshold defined by the noise level of the data gathering process.
-However, there has been no similar study on how instruction-tuning, the method that has enabled powerful LLMs such as ChatGPT, affects alignment to the human language system.
-In this work, we explore the impact of instruction-tuning on the alignment between LLMs and the human language system, considering two aspects: (1) brain alignment, which assesses how closely LLMs’ internal representations match neural activity patterns in the human language system, and (2) behavioral alignment, which evaluates the similarity between LLM behavior and human behavior.
-To conduct this study, both LLMs and human participants are presented with the same language stimuli comprised of naturalistic stories and sentences.
-For LLMs, we analyze their internal representations and per-word perplexity, while for humans, we use previously collected brain activity data from functional magnetic resonance imaging (fMRI) experiments and per-word reading times.
+Instruction-tuning is a widely adopted method that finetunes large language models (LLMs) on datasets containing many examples of different tasks and their descriptive instructions, enhancing their ability to generalize to previously unseen tasks by learning to follow provided instructions (Wang et al., 2022c).
+Despite costing only a small fraction of compute relative to pretraining (Chung et al., 2022), instruction-tuning yields large performance improvements on reasoning and problem-solving benchmarks.
+This generalization has allowed LLMs to tackle open-world reasoning tasks previously achievable only by humans (Zhang et al., 2023) while only using a few (or zero) task-specific training examples.
+In addition to teaching LLMs to understand and follow human instructions, instruction-tuning also improves the ability of LLMs to mimic human-written ground-truth outputs.
+This fluency allows them to produce more controllable and predictable output that is deemed (1) more desirable to human evaluators (Zhang et al., 2023; Chung et al., 2022; Wang et al., 2022b), (2) more aligned to human values (Chia et al., 2023), and (3) more stylistically similar to human outputs (Dasgupta et al., 2022; Safdari et al., 2023).
+Consequently, instruction-tuning yields LLMs more similar to humans in both capability and output resemblance.
+From a neuroscience perspective, this begs the question: Does instruction-tuning make LLMs more functionally similar to the human language system?
+Previous work has shown that models with higher task performance are more aligned to the human language system (Schrimpf et al., 2021; Goldstein et al., 2022; Caucheteux & King, 2022), hitting the estimated noise ceiling [^1] In fMRI recordings, a noise ceiling for representational similarity can be computed by sampling from the same participant twice, obtaining an upper limit for how well an ideal model could perform, defined by the internal consistency of neural responses and noise level of the data gathering process.
+on some datasets.
+However, there has been no similar study on how instruction-tuning affects alignment with the human language system.
+In this work, we measure the effect of instruction-tuning on the alignment between language mechanisms in LLMs and the human brain in two ways: (1) brain alignment, an “internal” metric that assesses how well internal feature representations of LLMs match neural representations in the human language system, and (2) behavioral alignment, an “external” metric which evaluates the similarity between LLM and human behavioral measurements.
+In our study, both LLMs and humans are presented with the same language stimuli comprised of naturalistic stories and sentences.
+For LLMs, we record their internal representations and per-word perplexity.
+For humans, we use previously recorded brain activity data from functional magnetic resonance imaging (fMRI) experiments and per-word reading times.
 To measure brain alignment, we use the Brain-Score (Schrimpf et al., 2018) linear predictivity metric, assessing how well LLM representations predict human brain activity in response to the same language stimuli (Jain & Huth, 2018; Toneva & Wehbe, 2019; Schrimpf et al., 2021; Oota et al., 2023), using data from three neural datasets: Pereira et al.
@@
 (2018) dataset.
-Perplexity for LLMs and reading times for humans offer insights into comprehension difficulty (Ehrlich & Rayner, 1981; Hale, 2001; Smith & Levy, 2013), allowing us to examine whether LLMs and humans share similarities in terms of which words and sentences they find challenging or surprising.
-Because models vary in their brain and behavioral alignment across different architectures and training objectives (Schrimpf et al., 2021), we estimate the effect of instruction-tuning across 17 instruction-tuned LLMs and 8 vanilla LLMs, and report a significant increase in brain alignment by instruction-tuned models compared to vanilla ones.
-To investigate why instruction-tuning increases alignment to human brain activity, we then estimate the contribution of various LLM properties towards brain alignment.
+Perplexity for LLMs and reading times for humans offer insights into comprehension difficulty (Ehrlich & Rayner, 1981; Hale, 2001; Smith & Levy, 2013), allowing us to examine whether LLMs match humans in their patterns of words and sentences they find challenging or surprising.
+Because models vary in their brain and behavioral alignment across different architectures and training objectives (Schrimpf et al., 2021), we estimate the effect of instruction-tuning by evaluating 8 vanilla LLMs and 17 LLMs that were further instruction-tuned, and report a significant increase in brain alignment due to instruction-tuning.
+To investigate why instruction-tuning increases alignment to human brain activity, we study the relationships between brain alignment and various LLM properties.
 Specifically, we compute Pearson correlations between an LLM’s brain alignment and its properties, including next-word prediction (NWP) ability, model size, a range of problem-solving abilities, and world knowledge spanning different domains.
-The evaluation of the latter two properties is based on the Big-Bench Hard benchmark (BBH) (Suzgun et al., 2022) and the Massive Multi-task Language Understanding benchmark (MMLU) (Hendrycks et al., 2021), respectively.
+We evaluated the latter two properties using the Big-Bench Hard benchmark (BBH; Suzgun et al., 2022) and the Massive Multi-task Language Understanding benchmark (MMLU; Hendrycks et al., 2021), respectively.
 We report three major findings:
   1. 1.
-    Instruction-tuning generally improves the alignment of LLM representations to brain activity, increasing brain alignment by 6.26.2% on average for the LLMs and neural datasets we tested (Figure 1).
+    Instruction-tuning generally aligns LLM representations to human brain activity, increasing brain alignment by $6.2$% on average (Figure 1).
   2. 2.
-    Investigating the factors underlying LLM-brain alignment, we find that world knowledge and model size are strongly correlated with brain alignment (r = $0.81$ and $0.95$ for instruction-tuned models, respectively; Figure 2).
+    Investigating the factors underlying LLM-brain alignment, we find that brain alignment is strongly correlated with world knowledge (r = $0.81$) and model size (r = $0.95$) (Figure 2).
   3. 3.
-    Surprisingly, our results generally indicate that instruction-tuning LLMs does not enhance behavioral alignment with human reading times.
-Furthermore, behavioral alignment on this dataset demonstrates poor correlations with all other measures we investigate, including task performance and model size (Figure 3).
+    Surprisingly, our results indicate that instruction-tuning LLMs generally does not enhance behavioral alignment with human reading times.
+Furthermore, behavioral alignment on this dataset is poorly correlated with all other measures we investigate, including task performance and model size (Figure 3).
 Figure 1: Instruction-tuning aligns LLM representations to human brain activity.
-(A) We evaluate brain alignment as the similarity of an LLM’s internal representations to human brain activity, using a linear predictivity metric in Brain-Score.
+(A) The same language stimuli are presented to LLMs and human participants.
+Next, we fit a linear regression model from LLM layer activations to fMRI responses in the human language system.
+We apply this linear model to predict held-out fMRI responses from the original corpus of recordings, and compute brain alignment as the Pearson correlation between the predicted and actual fMRI responses.
 We evaluate 25 vanilla and instruction-tuned LLMs with sizes between 77M and 33B parameters.
-We use 3 neural datasets of humans reading naturalistic stories and sentences: Pereira2018, Blank2014, and Wehbe2014, and compute the average across the three datasets.
+We compute the average across 3 neural datasets of humans reading naturalistic stories and sentences: Pereira2018, Blank2014, and Wehbe2014.
 (B) Instruction-tuning improves average brain alignment by 6.2% on average.
-Each point above the identity line represents an instruction-tuned LLM that has greater brain alignment than its vanilla counterpart.
-Error bars (here and elsewhere) represent median absolute deviation over human participants.
-(C) Instruction-tuning generally improves brain alignment on all three neural datasets.
-(D) We instruction-tune LLaMA-7B using the Alpaca dataset (“Instruction” model).
-We also train an ablation model with the same process and training data, but remove the instruction portion from each training sample (“No Instruction” model).
-This experiment shows that brain alignment improvements are due to both (1) training data (present in both models) and (2) the process of training LLMs to understand and follow instructions (present only in original model).
+Each point above the identity line is an instruction-tuned LLM with greater brain alignment than its vanilla version.
+Error bars, here and elsewhere, represent median absolute deviation over human participants.
+(C) Instruction-tuning improves average brain alignment on all three datasets.
+(D) We instruction-tune LLaMA-7B on the Alpaca dataset (“Instruction” model) and train an ablation model with the same data, but without the instruction in each training sample (“No Instruction” model).
+Our results show that brain alignment improvements are due to both (1) training data (present in both models) and (2) training LLMs to understand and follow instructions (present only in the first model).
 2 Background & Related Work
@@
 LM properties linked to Brain alignment.
-There is a growing body of work disentangling the contribution of various LM properties towards brain alignment.
+One exciting research area focuses on disentangling the contribution of various LM properties towards brain alignment.
 These include studying how brain alignment is driven by next-word prediction ability (Schrimpf et al., 2021; Caucheteux & King, 2022), multi-word semantics (Merlin & Toneva, 2022), performance on various NLP tasks (Oota et al., 2022), and model size (Antonello et al., 2023).
-To disentangle the contribution of various LM properties toward brain alignment, we use instruction-tuned LLMs.
+To contribute to this growing body of work, we use instruction-tuned LLMs.
 They are especially useful as they have been trained to respond to a standard question-answer format, allowing us to evaluate LLMs on a wide array of tasks and in a more fine-grained manner.
-We expand this area of research by identifying that world knowledge is a key property underlying LLM-brain alignment.
+We expand this area of research by demonstrating that world knowledge is a key property underlying LLM-brain alignment.
 3 Language Models
-We evaluate the brain alignment of 25 large language models (LLMs) from two model families: T5 (Raffel et al., 2020) and LLaMa (Touvron et al., 2023).
-T5 models are encoder-decoder LLMs pre-trained on the Colossal Common Crawl Corpus (C4), a corpus of 356 billion tokens, using a masked infilling objective, and then further finetuned on multi-task mixture of unsupervised and supervised tasks converted into a text-to-text format.
-In our study, we use all five T5 models with sizes between 77M to 11B parameters.
-LLaMA models (Touvron et al., 2023) are decoder-only LLMs trained on 1.6 trillion tokens from a mixture of corpora including C4, English CommonCrawl, Wikipedia, Github, and more.
-For LLaMA, we use the 7B, 13B, and 33B parameter versions in our study.
-For the instruction-tuned variants of T5 models, we utilize a variety of models finetuned on the FLAN suite (1515M examples for 1,836 different tasks accompanied by instructions, Chung et al., 2022), Alpaca (5252K instruction-following examples generated through methods inspired by Self-Instruct, Wang et al.
-(2022a), Taori et al., 2023), and GPT4ALL (437437K instruction-following examples generated with GPT-3.5-turbo, Anand et al., 2023) datasets.
-As for the LLaMa model family, we employ Vicuna’s 77B, 1313B, and 3333B models (Chiang et al., 2023), which are finetuned on user-shared conversations.
-Additionally, we incorporate the StableVicuna-13B model, which further refines the Vicuna-13B model using reinforcement learning from human feedback (RLHF) (Ouyang et al., 2022) on a range of conversational and instructional datasets.
-We also use the 77B version of Alpaca (Taori et al., 2023).
-Additional details about these LLMs can be found in Appendix A.
-Table 1: Example of instruction-tuning training data format: (Instruction, Input, Output) from the Alpaca dataset (Taori et al., 2023). The input field is optional for certain types of instructions.
+We evaluate the brain alignment of 25 large language models (LLMs) from two model families: T5 (Raffel et al., 2020) and LLaMa (Touvron et al., 2023a).
+T5 models are encoder-decoder LLMs pre-trained using a masked infilling objective on the Colossal Common Crawl Corpus (C4), a corpus of 356 billion tokens, and then further finetuned on a multi-task mixture of unsupervised and supervised tasks converted into a text-to-text format.
+We use all five T5 models, with sizes between 77M to 11B.
+LLaMA models (Touvron et al., 2023a) are decoder-only LLMs trained on 1.6 trillion tokens from a mixture of corpora including C4, English CommonCrawl, Wikipedia, GitHub.
+We use the 7B, 13B, and 33B versions.
+For the instruction-tuned variants of T5 models, we use a variety of models finetuned on FLAN ($15$M examples for 1,836 different tasks accompanied by instructions, Chung et al., 2022), Alpaca ($52$K instruction-following examples generated with methods inspired by Self-Instruct, Wang et al.
+(2022a), Taori et al., 2023), and GPT4ALL ($437$K instruction-following examples generated with GPT-3.5-turbo, Anand et al., 2023) datasets.
+For the instruction-tuned variants of LLaMa, we use Vicuna’s $7$B, $13$B, and $33$B models (Chiang et al., 2023), which are finetuned on user-shared conversations.
+We also use StableVicuna-13B, which further refines Vicuna-13B using reinforcement learning from human feedback (RLHF) (Ouyang et al., 2022) on a range of conversational and instructional datasets.
+We also use the $7$B version of Alpaca (Taori et al., 2023).
+We provide more details in Appendix A.
+We also tested 8 models from the GPT2 family; four vanilla models (Small, Medium, Large, XL) and their corresponding versions further instruction-tuned on Alpaca.
+However, the models performed close to random on reasoning and world knowledge benchmarks (BBH, MMLU).
+We hypothesize this is due to their size (all $<$ 1.5B parameters), making fine-tuning on a relatively small instruction-following dataset insufficient.
+Moreover, GPT2 was not originally pretrained to perform instruction-following tasks (unlike T5 models).
+We include their MMLU, BBH, brain and behavioral alignment results in Appendix G, H, and L.
 Instruction Input Output
-“Write a short paragraph “The importance of “The use of renewable energy is growing
-about the given topic.” using renewable energy.” rapidly in relevance and …”
+“Write a short paragraph “The importance of “The use of renewable energy
+about the given topic.” using renewable energy.” is growing…” (paragraph)
+Table 1: Example of Instruction-tuning training data format: (Instruction, Input, Output) from the Alpaca dataset (Taori et al., 2023). The input field is optional for certain instructions.
 4 Brain Alignment
-Brain alignment refers to the method of evaluating the similarity between LLM representations and human brain activity (Figure 1).
+Brain alignment refers to the method of evaluating the similarity between LLM representations and human brain activity (Figure 1A and Appendix C).
 This relies on fMRI recordings of human subjects while they read language stimuli on potentially any topic (here: Pereira et al., 2018; Blank et al., 2014; Wehbe et al., 2014).
 The same language stimuli from prior brain recordings are provided as input to LLMs, whose intermediate layer activations are extracted as their representations of the language stimuli.
-We follow a general approach previously used in several works (Schrimpf et al., 2018; 2021; Jain & Huth, 2018; Toneva & Wehbe, 2019; Oota et al., 2023; Aw & Toneva, 2023).
-Specifically, we use the linear predictivity metric implemented in Brain-Score (Schrimpf et al., 2020, Figure 1), first training a linear function to predict fMRI voxels associated with the human language system using LLM representations as input features.
-We then apply this linear function to held-out brain activity data from the original corpus of recordings, and evaluate the brain alignment of the LLM as the Pearson correlation between the predicted and actual brain activity data.
+We follow a general approach previously used in several works (Schrimpf et al., 2018; Jain & Huth, 2018; Toneva & Wehbe, 2019; Schrimpf et al., 2021; Oota et al., 2023; Aw & Toneva, 2023).
+Specifically, we use the linear predictivity metric implemented in Brain-Score (Schrimpf et al., 2018), which fits a linear regression model from LLM layer activations to fMRI responses in the human language system.
+We then apply this linear model to predict held-out fMRI responses from the original corpus of recordings, and compute brain alignment as the Pearson correlation between the predicted and actual fMRI responses.
 For each LLM, we evaluate its brain alignment for every layer (e.g., LLaMA-7B has 32 layers), and use the highest value as the LLM’s brain alignment value, following Schrimpf et al.
 (2018).
-Datasets   We use three fMRI datasets to measure the brain alignment of LLMs. Each dataset involves a different set of human participants and uses a different set of language stimuli.
-Pereira2018 (experiments 2 and 3 from Pereira et al., 2018): In experiment 2, nine participants read 384 sentences organized into 96 text passages.
-In experiment 3, six participants read 243 sentences in 72 text passages.
-Each sentence was displayed for four seconds on a screen.
-Blank2014 (Blank et al., 2014): The data consists of fMRI recordings of 5 human participants listening to naturalistic stories from the Natural Stories Corpus (Futrell et al., 2018).
-Wehbe2014 (Wehbe et al., 2014): The data includes fMRI recordings of 8 human participants reading chapter 9 of the book Harry Potter and the Sorceror’s Stone (Rowling et al., 1998).
+Datasets   We use three fMRI datasets to measure the brain alignment of LLMs. Each dataset involves a different set of human participants and a different set of language stimuli.
+Pereira2018 (experiments 2 and 3 from Pereira et al., 2018): In experiment 2, 9 participants read 384 sentences taken from 96 text passages.
+In experiment 3, 6 participants read 243 sentences from 72 text passages.
+Each sentence was displayed for 4 seconds on a screen.
+Blank2014 (Blank et al., 2014): The data consists of fMRI recordings of 5 human participants listening to 8 naturalistic stories from the Natural Stories Corpus (Futrell et al., 2018).
+Wehbe2014 (Wehbe et al., 2014): The data includes fMRI recordings of 8 human participants reading chapter 9 of the book Harry Potter and the Sorcerer’s Stone (Rowling et al., 1998).
 Participants read the chapter at a fixed interval of one word every 0.5 seconds.
@@
 We compute each LLM’s average brain alignment as the mean of its brain alignment on the 3 neural datasets.
-We find that instruction-tuning improves alignment by an average of 6.2% across all tested LLMs (Figure 1B).
-This holds across all three neural datasets, with average improvements of +6.9% on Pereira2018, +8.0% improvement on Blank2014, and +3.8% on Wehbe2014 (Figure 1C).
-This provides a strong signal that instruction-tuning improves brain alignment.
-Moreover, a smaller instruction-tuned model can attain higher brain alignment than a larger model from the same family that has not been instruction-tuned (e.g., Alpaca-7B v.s.
-LLaMa-13B, see detailed results in Appendix E).
-To longitudinally study how instruction-tuning aligns LLM representations to brain activity, we separately instruction-tune a LLaMA-7B model on the Stanford Alpaca instruction dataset (Taori et al., 2023) for 3 epochs.
-By evaluating checkpoints regularly during training, we find that instruction-tuning progressively improves brain alignment (Figure 1D).
-We also perform an ablation study to disambiguate the effect on brain alignment of (1) the instruction-following ability provided by the instruction-tuning step and (2) added training data.
-We fine-tune LLaMA-7B with the same process and data, but remove the instruction portion from each training sample.
-We observe that brain alignment of this ablated model increases during fine-tuning but stays lower than its instruction-following counterpart (Figure 1D).
-This shows that brain alignment improvements are due to both factors.
+We find that instruction-tuning improves brain alignment by an average of 6.2% across tested LLMs (Figure 1B).
+This holds across all three neural datasets, with improvements of +6.9% on Pereira2018, +8.0% improvement on Blank2014, and +3.8% on Wehbe2014 (Figure 1C).
+Moreover, a smaller instruction-tuned model can attain higher brain alignment than a larger vanilla model from the same family (e.g., Alpaca-7B v.s.
+LLaMa-13B, see Appendix G).
+However, some of the larger models (T5-XL, LLaMA) have minimal or no increases, possibly because the vanilla models are already close to the noise ceiling (Appendix G).
+Next, to longitudinally study how instruction-tuning aligns LLM representations to brain activity, we separately instruction-tune a LLaMA-7B model on the Stanford Alpaca instruction dataset (Taori et al., 2023) for 3 epochs.
+By evaluating checkpoints regularly during training, we find that instruction-tuning progressively improves brain alignment (Figure 1D, Appendix K).
+We also disambiguate the effect on brain alignment of (1) the instruction-following ability gained from instruction-tuning and (2) added training data.
+We fine-tune LLaMA-7B with the same process and data, but remove the instruction from each training sample.
+Brain alignment of this ablated model increases during fine-tuning but stays lower than its instruction-following counterpart (Figure 1D), indicating that brain alignment improvements are due to both factors.
+In Appendix J, we also report preliminary results using recent models from the Gemma (Team et al., 2024) and LLaMA-2 (Touvron et al., 2023b) families and observe similar trends.
+Further, these trends hold across different similarity metrics besides Linear predictivity, e.g., RSA and CKA (Appendix J).
 Figure 2: World knowledge and model size are important factors underlying LLM-brain alignment.
 To identify factors underlying brain alignment, we test Pearson correlations between brain alignment and various LLM properties, such as model size, world knowledge in various domains (MMLU benchmark), and various types of problem-solving abilities (BBH benchmark).
-Insets display results on individual datasets, with stars reflecting statistical significance (n.s.
-= p >> 0.05, * = p << 0.05, ** = p << 0.005, etc.) (A) Brain alignment is significantly and strongly correlated with world knowledge as evaluated by the MMLU Overall score (r = 0.81), which reports the mean performance across all MMLU subjects.
+In the figure, insets display results on individual datasets, with stars reflecting statistical significance (n.s.
+= p $>$ 0.05, * = p $<$ 0.05, ** = p $<$ 0.005, etc.) (A) Brain alignment is significantly and strongly correlated with world knowledge as evaluated by the MMLU Overall score (r = 0.81), which reports the mean performance across all MMLU subjects.
 (B) Brain alignment is significantly and strongly correlated with performance on the world knowledge task category in BBH (r = 0.68).
 (C) Brain alignment is significantly and strongly correlated with model size (logarithm of number of parameters) (r = 0.95).
+In Appendix I, we provide a larger version of this figure with labels for each data point.
+Benchmark, Task category Correlation (r) to Adjusted No. Avg. Model
+Brain Alignment p-value tasks Accuracy
+MMLU, Overall Score 0.809 $<$ 0.0005 57 0.36
+MMLU, STEM 0.792 $<$ 0.0005 18 0.28
+MMLU, Humanities 0.791 $<$ 0.0005 13 0.34
+MMLU, Social Sciences 0.807 $<$ 0.0005 12 0.41
+MMLU, Others 0.809 $<$ 0.0005 14 0.40
+BBH, Overall score 0.384 0.18 23 0.28
+BBH, Algorithmic reasoning 0.194 0.56 8 0.22
+BBH, Language understanding 0.163 0.59 3 0.43
+BBH, World knowledge 0.679 $<$ 0.005 5 0.36
+BBH, Multilingual reasoning -0.035 0.90 1 0.19
+BBH, Others 0.478 0.08 6 0.27
+Table 2: Brain alignment strongly correlates with world knowledge in all subject domains in MMLU and the world knowledge category in BBH.
+Brain alignment is not significantly correlated with other problem-solving abilities in BBH (e.g., algorithmic or multilingual reasoning).
+We obtain p-values after performing false discovery rate (FDR) correction.
 4.2 Factors underlying LLM-brain alignment
-To identify factors underlying the representational similarity between LLMs and human brains, we compute the Pearson correlation between LLM brain alignment and various properties of LLMs: performance on a benchmark involving different reasoning abilities (BBH benchmark; Suzgun et al., 2022), performance on a benchmark requiring domain-specific world knowledge (MMLU; Hendrycks et al., 2021), language modeling ability, and model size.
-World Knowledge and Reasoning abilities   MMLU is designed to measure the world knowledge of LLMs across many subject domains.
-It contains 57 tasks, categorized by the subject domain of world knowledge tested: STEM, Humanities, Social Sciences, and Others (a broad category ranging from finance to marketing to professional medicine).
-BBH contains 23 tasks, grouped into four categories based on the type of problem-solving ability tested: Algorithmic and Multi-Step Arithmetic Reasoning; Natural Language Understanding; Use of World Knowledge; and Multilingual Knowledge and Reasoning.
-For both benchmarks, we use category classifications from the original papers.
-We measure the performance of LLMs on BBH and MMLU using the instruct-eval repository [^2] https://github.com/declare-lab/instruct-eval with default settings (3-shots, 5-shots respectively) and preset prompts.
-We measure the Pearson correlation (and its p-value) between LLM brain alignment and performance on each category of the MMLU and BBH benchmarks.
-We obtain p-values after performing false discovery rate correction, ensuring our results are significant and not due to random chance.
+To study why LLMs and human brains align in their representations, we compute the Pearson correlation between LLM brain alignment and various properties of LLMs: performance on a benchmark testing various reasoning abilities (BBH; Suzgun et al., 2022), performance on a benchmark testing world knowledge in various domains (MMLU; Hendrycks et al., 2021), language modeling ability, and model size.
+MMLU and BBH.
+MMLU is designed to measure the world knowledge of LLMs across many subject domains.
+It contains 57 tasks, categorized into four world knowledge subject domains: STEM, Humanities, Social Sciences, and Others (a broad category ranging from finance to marketing to professional medicine).
+BBH contains 23 tasks, grouped into four types of problem-solving: Algorithmic and Multi-Step Arithmetic Reasoning; Natural Language Understanding; Use of World Knowledge; and Multilingual Knowledge and Reasoning.
+For both benchmarks, we follow the category classifications from the original papers.
+We perform the evaluations using the instruct-eval repository [^2] https://github.com/declare-lab/instruct-eval with default settings (3-shots for BBH, 5-shots for MMLU) and preset prompts.
+We measure the Pearson correlation (and its p-value) between LLM-brain alignment and performance in each category of MMLU and BBH.
+We obtain p-values after false discovery rate (FDR) correction.
+World Knowledge.
 We find that brain alignment is significantly and strongly correlated with world knowledge.
-On the MMLU benchmark, we observe a high correlation between brain alignment scores and the MMLU Overall score (r = 0.81), which reports the mean performance across all world knowledge subject domains on MMLU (Figure 2A).
-Similarly, brain alignment is also significantly and strongly correlated with the mean performance on tasks included in the world knowledge category of the BBH benchmark (r = 0.68; Figure 2B).
-Interestingly, we do not find strong correlations with other dimensions of the BBH benchmark (e.g., Algorithmic reasoning and Multilingual reasoning, see Table 2), though this could also be due to limitations of the tested models (most are primarily pretrained on English language data), as indicated by their low raw performance scores on some tasks.
+Brain alignment is highly correlated with the MMLU Overall score (r = 0.81, p $<$ 0.001, Figure 2A), which reports the mean performance across all world knowledge subject domains on MMLU.
+Similarly, brain alignment is also strongly correlated with performance on tasks in the world knowledge category of BBH (r = 0.68, p $<$ 0.005; Figure 2B).
+Notably, brain alignment is poorly correlated with all other dimensions of BBH (see Table 2), though this could also be due to limitations of the tested models, as indicated by their low raw performance scores on some tasks.
 Overall, our results provide a strong signal that more accessible representations of world knowledge are a key factor in aligning LLM representations to human brain activity.
-Language Modeling Ability   Prior works have shown correlations between brain alignment and next-word prediction (NWP) ability (Caucheteux & King, 2022; Schrimpf et al., 2021).
-We find similar results for correlation between brain alignment and NWP loss (r = -0.54, Appendix G).
-Interestingly, the strength of correlation between brain alignment and world knowledge performance (r = 0.81) is greater than that between brain alignment and NWP loss (r = -0.54).
-This result suggests that world knowledge understanding is a better predictor of brain alignment than NWP ability.
-Model Size   Finally, we find that brain alignment is significantly and strongly correlated with model size (r = 0.95), as measured by the logarithm of the number of model parameters (Figure 2C).
+Language Modeling Ability.
+Prior works have shown correlations between brain alignment and next-word prediction (NWP) ability (Caucheteux & King, 2022; Schrimpf et al., 2021).
+We find similar results for correlation between brain alignment and NWP loss (r = -0.54, p $<$ 0.05, Appendix I).
+Interestingly, the strength of the correlation is weaker than that between brain alignment and world knowledge performance (r = 0.81).
+This suggests that world knowledge understanding is a better predictor of brain alignment than NWP ability.
+Model Size.
+Finally, we find that brain alignment is significantly and strongly correlated with model size (r = 0.95, p $<$ 0.001, Figure 2C), as measured by the logarithm of the number of model parameters.
 Schrimpf et al.
@@
 For example, LLaMA-13B obtains brain alignment of 0.220, Vicuna-13B obtains 0.229, LLaMA-33B obtains 0.227, and Vicuna-33B obtains 0.232.
-Hence, Vicuna-13B has greater brain alignment than LLaMA-33B despite being less than 40% its size.
+Hence, Vicuna-13B has greater brain alignment than LLaMA-33B, due to instruction-tuning, despite being less than 40% its size.
 We observe a similar trend in another four models: T5-base, Flan-T5-base, T5-large, Flan-T5-large.
-Also, prior works have shown that large random models achieve poor brain alignment (Schrimpf et al., 2021).
-These results demonstrate that there are LLM properties aside from model size that contribute significantly to brain alignment.
-Datasets   We observe significant correlations between brain alignment and world knowledge and model size on Pereira2018 and Wehbe2014, but surprisingly not Blank2014.
-We hypothesize that it may be because Blank2014 has a small participant size (N = 5) that leads to greater noise in the results, compared to Pereira2018 and Wehbe2014.
-Table 2: Brain alignment strongly correlates with world knowledge across all subject domains in MMLU, as well as the world knowledge category in BBH.
-At the same time, brain alignment is not significantly correlated with all other types of problem-solving abilities in BBH (e.g., algorithmic or multilingual reasoning).
-We obtain p-values after performing false discovery rate correction, ensuring our results are significant and not due to random chance.
-Task category Brain Alignment corrected Number Average Model
-Correlation (r) p-value of tasks Performance
-MMLU – Overall Score 0.809 0.000329 57 0.36
-MMLU – STEM 0.792 0.000343 18 0.28
-MMLU – Humanities 0.791 0.000343 13 0.34
-MMLU – Social Sciences 0.807 0.000329 12 0.41
-MMLU – Others 0.809 0.000329 14 0.40
-BBH – Overall score 0.384 0.177 23 0.28
-BBH – Algorithmic reasoning 0.194 0.558 8 0.22
-BBH – Language understanding 0.163 0.585 3 0.43
-BBH – World knowledge 0.679 0.005 5 0.36
-BBH – Multilingual reasoning -0.035 0.895 1 0.19
-BBH – Others 0.478 0.083 6 0.27
+Also, prior works showed that large random models achieve poor brain alignment (Schrimpf et al., 2021).
+These results demonstrate there are LLM properties aside from model size that contribute significantly to brain alignment.
+Neural datasets.
+Brain alignment is strongly correlated with world knowledge and model size on Pereira2018 and Wehbe2014, but less strongly correlated on Blank2014, possibly because Blank2014 has fewer participants (N = 5) leading to greater noise in the results.
 5 Behavioral Alignment
@@
 Furthermore, behavioral alignment correlates poorly with all other tested measures: world knowledge, model size, and next-word prediction (NWP) ability.
-(A) We present the same language stimuli to LLMs and human participants, using the Futrell2018 benchmark in Brain-Score, which contains naturalistic stories.
-We compute the behavioral alignment as the Pearson correlation between LLM perplexity for each word and human reading times for each word.
-(B) Instruction-tuning does not generally improve behavioral alignment.
-Furthermore, behavioral alignment is poorly and not significantly correlated with all other measures: (C) world knowledge on MMLU (r = 0.08, p = 0.76), (D) model size (r = 0.26, p = 0.31), (E) NWP loss for T5 models (r = -0.2, p = 0.54), and (F) NWP loss for LLaMA models (r = 0.68, p = 0.21).
-In the previous section, we show that instruction-tuning aligns the internal representations of LLMs to human brain recordings (Section 4.1).
+To compute behavioral alignment, we use the Futrell2018 benchmark in Brain-Score.
+The same language stimuli (naturalistic stories) are presented to LLMs and human participants.
+We then compute the Pearson correlation between per-word LLM perplexity and human reading times as the behavioral alignment.
+(A) Instruction-tuning does not generally improve behavioral alignment.
+Furthermore, behavioral alignment is poorly and not significantly correlated with all other measures: (B) world knowledge (p = 0.76), (C) model size (p = 0.31), (D) NWP loss for T5 models (p = 0.54), and (E) NWP loss for LLaMA models (p = 0.21).
+In Appendix M, we provide a larger, labeled version of this figure.
+In the previous section, we show that instruction-tuning aligns the internal representations of LLMs to human brain activity (Section 4.1).
 In this section, we explore whether instruction-tuning also aligns LLM behavior to human behavior.
-Following the approach previously proposed by Schrimpf et al.
-(2021) and implemented in the Brain-Score package (Schrimpf et al., 2020), we measure behavioral alignment by evaluating the similarity between LLM per-word perplexity and human per-word reading times, given the same language stimuli (Figure 3A).
+Following the approach of Schrimpf et al.
+(2021) implemented in the Brain-Score package (Schrimpf et al., 2020), we measure behavioral alignment by evaluating the similarity between LLM per-word perplexity and human per-word reading times given the same language stimuli (Appendix D).
 We use the self-paced reading times dataset from Futrell et al.
 (2018), consisting of the reading times of 179 human participants recorded while they were visually presented with 10 naturalistic stories.
-We provide language stimuli from this data as input to LLMs and measure their per-word perplexity.
-Finally, we evaluate behavioral alignment by computing the Pearson correlation between per-word LLM perplexity and per-word human reading times.
-Using the same set of models as in the previous section, we compare the behavioral alignment of each instruction-tuned LLM against its vanilla counterpart.
-Our results generally indicate that instruction-tuning LLMs does not improve behavioral alignment to human reading times (Figure 3B).
+We provide language stimuli from this data as input to LLMs and evaluate behavioral alignment by computing the Pearson correlation between per-word LLM perplexity and per-word human reading times.
+One potential concern in the behavioral alignment methodology is that human reading time is influenced by word length; longer words may take more time to read, even if they are not more challenging or surprising.
+Following Schrimpf et al.
+(2021), we address this by computing the surprisal of a multi-token word as the sum of the surprisals of its individual tokens.
+Importantly, we use consistent tokenizers when comparing vanilla LLMs against their instruction-tuned versions, thereby controlling for the effect of word length when evaluating the impact of instruction-tuning on behavioral alignment.
+5.1 Instruction-tuning generally does not improve behavioral alignment
+Using the same models as in Section 4, we compare the behavioral alignment of each instruction-tuned LLM against its vanilla counterpart.
+Our results indicate instruction-tuning generally does not improve behavioral alignment to human reading times (Figure 3A).
 For half of the LLMs tested, it results in no change or reduced behavioral alignment.
-Then, we test the correlation between LLM behavioral alignment and model size, next-word prediction ability, various reasoning abilities (as measured by performance on the BBH benchmark), and world knowledge across various domains (as measured by performance on the MMLU benchmark).
-Contrary to our findings on the correlation between brain alignment and model size and world knowledge (Section 4.2), we do not find that these factors are correlated with the behavioral alignment of models: world knowledge in Figure 3C (r = 0.08, p = 0.76), model size in Figure 3D (r = 0.26, p = 0.31), next-word prediction loss for T5 models in Figure 3E (r = -0.2, p = 0.54), and next-word prediction loss for LLaMA models in Figure 3F (r = 0.68, p = 0.21).
-We discuss our interpretations of these results and possible explanations in Section 6.2.
+Our results align with recent studies (Gao et al., 2023; Steuer et al., 2023) that show improving LLM performance does not necessarily improve alignment with human reading behavior.
+5.2 Factors underlying behavioral alignment
+Next, we test the correlation between LLM behavioral alignment and model size, next-word prediction ability, various reasoning abilities (measured by BBH), and world knowledge across various domains (measured by MMLU).
+Contrary to our findings on the brain alignment correlations with model size and world knowledge (Section 4.2), we find that LLM behavioral alignment is not correlated with these factors: world knowledge (p = 0.76, Figure 3B), model size (p = 0.31, Figure 3C), next-word prediction loss for T5 models (p = 0.54, Figure 3D), and next-word prediction loss for LLaMA models (p = 0.21, Figure 3E).
 6 Discussion
 6.1 Implications for NLP: Building LLMs
-Using brain alignment to interpret LLMs.
-Representations in the human brain support a wide range of downstream tasks that current models still struggle on.
-Prior works have used brain activity results to interpret neural networks (Dong & Toneva, 2023) and build more performant models (Dapello et al., 2020; 2022; Safarani et al., 2021).
-Instruction-tuning has emerged as a breakthrough technique to improve LLM abilities, quality of outputs, and allow LLMs to adapt to new tasks with minimal task-specific training.
-However, the manner in which instruction-tuning alters the internal representations of LLMs to achieve these improvements remains an open question.
-Brain activity data provides a neuroscientific angle to investigate this question.
-Our results show that instruction-tuning LLMs improves both world knowledge representations and human brain alignment.
-They suggest that instruction-tuning improves the ability of an LLM to retrieve world knowledge stored in its model weights and that the mechanisms that encode world knowledge in LLMs also improve representational alignment to the human brain.
-Overall, our results highlight the importance of integrating world knowledge in the development of future LLMs.
+Using brain alignment for LLM interpretability.
+Prior works have used human brain activity to interpret models (Dong & Toneva, 2023) and build more performant models (Dapello et al., 2020; Safarani et al., 2021; Dapello et al., 2022).
+Instruction-tuning has emerged as a breakthrough technique to improve LLM abilities and quality of outputs, and allows LLMs to adapt to new tasks with minimal task-specific training.
+However, how instruction-tuning alters LLM internal representations to achieve these improvements remains an open question.
+Brain activity provides a neuroscientific angle to investigate this question.
+Our results show that instruction-tuning improves both performance on world knowledge benchmarks and human brain alignment, suggesting that the mechanisms that encode world knowledge in LLMs also align LLM representations to the human brain.
 6.2 Implications for Neuroscience: Studying LLM-Human Alignment
 Instruction-tuned LLMs are useful for studying LLM properties underlying brain and behavioral alignment.
-To identify why LLM and human brains exhibit representational similarities, prior work has mostly focused on high-level properties such as model size (Antonello et al., 2023), and external behaviors such as predicting missing words Schrimpf et al.
-(2021); Caucheteux & King (2022).
-However, a key to understanding these similarities is to identify lower-level or internal properties of LLMs that underlie brain alignment.
-This includes the amount of knowledge LLMs contain, e.g., factual (AlKhamissi et al., 2022) and commonsense (Sap et al., 2020; Bosselut et al., 2019).
-Our work shows that we can harness instruction-tuned LLMs for this purpose as they have been trained to respond to a general question format, allowing us to evaluate LLMs in a more fine-grained manner.
-This allows the study of both internal (e.g., knowledge) and external (e.g., behavior) properties of LLMs, and how they correlate with brain and behavioral alignment.
+To identify why LLM and human brains share representational similarities, prior work has mostly focused on high-level properties such as model size (Antonello et al., 2023), as well as external behaviors such as predicting missing words (Schrimpf et al., 2021; Caucheteux & King, 2022).
+However, a key to understanding these similarities is to identify internal properties of LLMs that underlie brain alignment, including the amount of knowledge LLMs encode, e.g., factual (AlKhamissi et al., 2022) and commonsense (Bosselut et al., 2019; Sap et al., 2020).
+Our work is the first to show that we can harness instruction-tuned LLMs for this purpose.
+Because they have been trained to respond to a general instruction format, we can evaluate LLMs on diverse tasks in a fine-grained manner, allowing the study of LLM properties both internal (e.g., knowledge) and external (e.g., behavior), and how they correlate with brain and behavioral alignment.
+World knowledge shapes language comprehension and brain activity.
+Our results show that world knowledge is a key factor in LLM brain alignment.
+LLMs demonstrating greater world knowledge across all tested subject domains have representations more similar to the human brain.
+This suggests that world knowledge influences human brain activity and shapes the language comprehension systems in our brain.
+6.3 Limitations and Future Work
 Examining more dimensions of behavior.
-To evaluate behavioral alignment, our work and many prior works compare LM and human next-word surprisal on reading tasks (Wilcox et al., 2020; Schrimpf et al., 2021; Eghbal A.
-Hosseini et al., 2023).
-This evaluates only a single dimension of LM and human behavior (per-word perplexity and reading times).
-On the models we test here, behavioral alignment is not significantly correlated with model size, world knowledge, or next-word prediction ability.
-While next-word prediction performance correlates with alignment to human reading times across a broad band of models (Schrimpf et al., 2021), this trend does not hold up in recent Transformer-based LMs (Oh & Schuler, 2023), having a surprising negative correlation with parameter count (Oh et al., 2022).
-Our results highlight the need to create more benchmarks to expand the dimensions of behavior examined for both LLMs and humans, in order to holistically evaluate LLM behavior, as well as LLM-human behavioral alignment.
+Our work and many prior works compare LLM and human next-word surprisal on reading tasks (Wilcox et al., 2020; Schrimpf et al., 2021; Eghbal A.
+Hosseini et al., 2023), which evaluates only a single behavioral dimension for LLMs (per-word perplexity) and humans (reading times).
+For the models we tested, behavioral alignment is not significantly correlated with model size, world knowledge, or next-word prediction ability.
+While next-word prediction performance correlates with alignment to human reading times across many LMs (Schrimpf et al., 2021), this trend does not hold up in recent transformer models (Oh & Schuler, 2023), having a surprising negative correlation with parameter count (Oh et al., 2022).
+In the future, we hope to evaluate LLMs on a wider range of behavioral dimensions (e.g., van Duijn et al., 2023; Koo et al., 2023; Kauf et al., 2024) to more holistically evaluate LLM-human behavioral alignment.
 Brain alignment datasets with humans performing diverse tasks.
-Our work studies brain alignment using neural datasets that are limited to humans reading naturalistic stories and sentences in English.
-Unfortunately, there does not exist brain activity data for human participants attempting the BBH and MMLU benchmarks.
-This may explain why brain alignment is not significantly correlated with many categories of problem-solving on BBH, e.g., language understanding.
+We study brain alignment using neural datasets of humans reading naturalistic stories and sentences in English.
+It would be interesting to study brain alignment to human participants attempting the BBH and MMLU benchmarks, but this data unfortunately does not exist.
+This may explain why brain alignment is not significantly correlated with many categories of problem-solving on BBH, e.g., algorithmic reasoning.
 In the future, we hope to study brain alignment with human participants performing more diverse sets of tasks, e.g., reading computer program code (Ivanova et al., 2020).
 This can identify more factors underlying LLM-brain alignment, and provide insights into how brain activity and the human language system may be shaped by various forms of problem-solving.
-Furthermore, for the neural datasets in our work, many of the larger models exceed the noise ceiling estimates (Appendix E), highlighting the need for more neural datasets (with better ways of computing noise ceiling estimates).
-World knowledge shapes brain activity.
-Our results show that world knowledge is a key factor in aligning LLM representations to human brain activity.
-LLMs with greater world knowledge across all tested subject domains produce representations that align more closely to human brain activity.
-Analogously, this suggests that world knowledge influences human brain activity, shaping the language comprehension systems in the brain.
-6.3 Limitations and Future Work
-Limited model families.
-A model’s performance on MMLU and BBH is affected not only by the quantity of knowledge it possesses, but also by its ability to follow each benchmark’s instruction format.
-Hence, we restricted our study to two model families (T5 and LLaMA).
-This attempts to ensure that the models are trained on a similar instruction format, so they have similar ability to follow the instruction format of MMLU and BBH.
-Thus, any differences in their MMLU or BBH performance would point to differences in the quantity of knowledge they possess.
-However, we hope to add a greater variety of LLMs in future work to strengthen the generalizability of our claims.
-Computational cost.
-We investigate 25 LLMs, with the largest models having 33B parameters, on many datasets: brain alignment (Pereira2018, Blank2014, Wehbe2014), behavioral alignment (Futrell2018), next-word prediction (WikiText-2) and other evaluations (BBH, MMLU).
-Unfortunately, this makes it computationally expensive to replicate our experiments.
-In the hopes of addressing this concern, we provide our full results in Appendices E, F, G, H, and I.
+Furthermore, some of the larger models exceed the noise ceiling estimates in our neural datasets (Appendix G), highlighting the need for more neural datasets and better ways of computing noise ceiling estimates.
 7 Conclusion
@@
 We find that instruction-tuning generally improves the alignment of LLM representations to brain activity.
-Delving into the factors underlying LLM-brain alignment, we discover that world knowledge and model size are key determinants of brain alignment.
-This suggests that world knowledge helps shape representations in the human language system, and highlights the significance of integrating world knowledge in the development of future LLMs.
-Reproducibility Statement
-All models and code repositories utilized in this study are open-source and their corresponding links are provided in Appendix B and D for full transparency and reproducibility.
-For the calculation of brain and behavioral alignment, we relied on the Brain-Score repository (www.github.com/brain-score/language), which serves as a publicly accessible resource for conducting these assessments.
+Exploring the factors underlying LLM-brain alignment, we discover that world knowledge and model size are key determinants of brain alignment.
+This suggests that world knowledge helps shape representations in the human language system, and highlights the importance of integrating world knowledge in developing future LLMs.
+8 Reproducibility Statement
+We provide our full results in Appendices G, H, and L.
+We consider this important as our experiments may be computationally expensive to replicate.
+We investigated 25 LLMs, with the largest models having 33B parameters, on many datasets: brain alignment (Pereira2018, Blank2014, Wehbe2014), behavioral alignment (Futrell2018), next-word prediction (WikiText-2) and other evaluations (BBH, MMLU).
+All models and code repositories used in our study are open-source and their corresponding links are provided in Appendix B and F for full transparency and reproducibility.
+To calculate brain and behavioral alignment, we used the Brain-Score repository (www.github.com/brain-score/language), a publicly accessible resource for conducting these assessments.
 We encourage researchers interested in replicating our findings to refer to the provided links and consult the Brain-Score repository for further details on datasets and the evaluation process.
-For measuring the performance of LLMs on BBH and MMLU benchmarks, we use the broadly-used instruct-eval repository (https://github.com/declare-lab/instruct-eval) with default settings.
+To evaluate LLMs on BBH and MMLU, we utilize the widely-used instruct-eval repository (https://github.com/declare-lab/instruct-eval) with default settings.
 Appendix A Language Models: Parameter count and Number of Layers
-Table 3: Parameter count and number of layers for all 25 vanilla and instruction-tuned LLMs.
-The upper part contains encoder-decoder models of the T5 family, the lower parts decoder-only models of the LLaMA family.
-For the parameter count, “M” refers to million and “B” refers to billion.
-The number of layers for T5 models is a sum of the number of encoder and decoder layers.
 Model Parameter Count Number of Layers
@@
 vicuna-33b 33 B 60
+gpt2-small 124 M 12
+gpt2-small-alpaca 124 M 12
+gpt2-medium 355 M 24
+gpt2-medium-alpaca 355 M 24
+gpt2-large 774 M 36
+gpt2-large-alpaca 774 M 36
+gpt2-xl 1.5 B 48
+gpt2-xl-alpaca 1.5 B 48
+Table 3: Parameter count and number of layers for all vanilla and instruction-tuned LLMs.
+For the parameter count, “M” refers to million and “B” refers to billion.
+The number of layers for T5 models is a sum of the number of encoder and decoder layers.
 Appendix B Language Models: Links to models weights
-Table 4: Link to model weights for all 25 vanilla and instruction-tuned LLMs.
-The upper part contains encoder-decoder models of the T5 family, the lower parts decoder-only models of the LLaMA family.
-We provide these links for reproducibility purposes.
 Model Link to model weights
@@
 vicuna-33b www.huggingface.co/lmsys/vicuna-33b-v1.3
-Appendix C MMLU and BBH benchmarks
+gpt2-small https://huggingface.co/openai-community/gpt2
+gpt2-small-alpaca https://huggingface.co/vicgalle/gpt2-alpaca
+gpt2-medium https://huggingface.co/openai-community/gpt2-medium
+gpt2-medium-alpaca https://huggingface.co/linkanjarad/GPT2-Medium-Alpaca-355m
+gpt2-large https://huggingface.co/openai-community/gpt2-large
+gpt2-large-alpaca https://huggingface.co/reasonwang/gpt2-large-alpaca
+gpt2-xl https://huggingface.co/openai-community/gpt2-xl
+gpt2-xl-alpaca https://huggingface.co/Rachneet/gpt2-xl-alpaca
+Table 4: Link to model weights for all vanilla and instruction-tuned LLMs. We provide these links for reproducibility purposes.
+Appendix C Method for computing Brain alignment
+Figure 4: Method for computing Brain alignment, the similarity of an LLM’s internal representations to human brain activity.
+We follow a general approach for aligning LLM representations to fMRI brain activity used in prior works (Jain & Huth, 2018; Toneva & Wehbe, 2019; Schrimpf et al., 2021).
+The approaches used for the three neural datasets are similar.
+Hence, we explain the main details only for Wehbe2014, followed by key differences for the other two datasets.
+C.1 Wehbe2014
+Wehbe2014 (Wehbe et al., 2014): The data includes fMRI recordings of 8 human participants reading chapter 9 of the book Harry Potter and the Sorceror’s Stone (Rowling et al., 1998).
+Participants read the chapter at a fixed interval of one word every 0.5 seconds.
+The chapter contains 5,176 words.
+LLM representations. We extract LLM representations of the text at each layer $\in\mathbb{R}^{w\times d}$, where $w$ is the number of words in the text and $d$ is the embedding size of the LLM.
+fMRI activity.
+Data analyses were performed on fMRI BOLD signals extracted from the language network for all three neural datasets.
+fMRI activity was recorded at regular intervals of 2 seconds, with a total of 1,351 recorded time-points (TRs, times of repetition).
+We construct a stimulus-response matrix for fMRI activity $\in\mathbb{R}^{TR\times v_{i}}$, where $TR$ is the number of TRs and $v_{i}$ is the number of voxels in the brain of participant $i$.
+Alignment of LLM representations to fMRI activity.
+We use PCA to reduce the dimensionality of the LLM representations to result in a matrix $\in\mathbb{R}^{w\times 10}$.
+As the number of words in the text is greater than the number of TRs, we down-sample the LLM word-level representations to the TR rate by averaging the LLM representations to the corresponding TRs, producing a matrix $\in\mathbb{R}^{TR\times 10}$.
+As the response recorded by fMRI peaks about 6 seconds after stimulus onset, we follow prior methods by including preceding time-points for each time-point.
+The final LLM representations are constructed by concatenating, for each TR, the LLM representations corresponding to the previous 4 TRs, producing a matrix $\in\mathbb{R}^{TR\times 40}$.
+This accounts for the lag in the hemodynamic response that fMRI records.
+Finally, we learn a linear function, regularized by the ridge penalty, that uses the modified LLM representations $\in\mathbb{R}^{TR\times 40}$ to predict the fMRI activity of participants $\in\mathbb{R}^{TR\times v_{i}}$.
+We train the function using 4-fold cross-validation, where each fold corresponds to a separate run of fMRI collection.
+We also remove 10 TRs from the LLM representations and fMRI activity from the ends of each test run to avoid train-test overlap.
+To compute the brain alignment, we evaluate the Pearson correlation (r) between the predicted fMRI activity $\in\mathbb{R}^{TR\times v_{i}}$ and recorded fMRI activity $\in\mathbb{R}^{TR\times v_{i}}$.
+C.2 Blank2014
+Blank2014 (Blank et al., 2014): The data consists of fMRI recordings of 5 human participants listening to 8 naturalistic stories from the Natural Stories Corpus (Futrell et al., 2018).
+fMRI activity was recorded at regular intervals of 2 seconds, with a total of 1,317 TRs.
+We average the BOLD signals across voxels within each language-responsive region of interest (ROI) of each participant to increase the signal-to-noise ratio, following Schrimpf et al.
+(2021).
+As there are 60 ROIs in total across the 5 participants, this produces a stimulus-response matrix $\in\mathbb{R}^{1,317\times 60}$.
+C.3 Pereira2018
+Pereira2018 (experiments 2 and 3 from Pereira et al., 2018): In experiment 2, 9 participants read 384 sentences taken from 96 text passages.
+In experiment 3, 6 participants read 243 sentences from 72 text passages.
+Each sentence was displayed for 4 seconds on a screen.
+We average the fMRI responses for each sentence, resulting in one data point per sentence per language-responsive voxel of each participant.
+For experiment 2, there are 384 sentences and 12,195 language-responsive voxels across the participants.
+For experiment 3, there are 243 sentences and 8,121 language-responsive voxels across the participants.
+We concatenate the fMRI responses across sentences and participants for each experiment.
+This produces a stimulus-response matrix $\in\mathbb{R}^{384\times 12,195}$ for experiment 2 and a stimulus-response matrix $\in\mathbb{R}^{243\times 8121}$ for experiment 3.
+Appendix D Method for computing Behavioral alignment
+Figure 5: Method for computing Behavioral alignment.
+The same language stimuli are presented to LLMs and human participants, using the Futrell2018 benchmark in Brain-Score, which contains naturalistic stories.
+We compute the behavioral alignment as the Pearson correlation between LLM perplexity for each word and human reading times for each word.
+Futrell2018 (Futrell et al., 2018): The data consists of reading times (RTs) of 179 participants on a total of 10,256 words across 485 sentences, taken from 10 stories from the Natural Stories Corpus (Futrell et al., 2018).
+The stories were presented online to Amazon Mechanical Turk users visually in a dashed moving window display.
+Participants press a button to reveal each consecutive word.
+When they press the button, the most recent word is converted back to dashes, and the next word is revealed.
+The time between button presses offers an estimate of comprehension difficulty (Ehrlich & Rayner, 1981; Hale, 2001; Smith & Levy, 2013).
+We construct a stimulus-response matrix of RTs $\in\mathbb{R}^{10,256\times 179}$.
+We provide the same stimuli to an LLM and record its perplexity for each word, producing an LLM perplexity matrix $\in\mathbb{R}^{10,256}$.
+We compute behavioral alignment as the Pearson correlation between per-word LLM perplexity and human RTs.
+Perplexity for LLMs and reading times for humans offer insights into comprehension difficulty (Ehrlich & Rayner, 1981; Hale, 2001; Smith & Levy, 2013), allowing us to examine whether LLMs and humans share similarities in terms of which words and sentences they find challenging or surprising.
+Appendix E MMLU and BBH benchmarks
 MMLU is designed to measure knowledge from many domains Hendrycks et al.
@@
 We measure the performance of LLMs on BBH and MMLU using the instruct-eval repository [^3] https://github.com/declare-lab/instruct-eval with default settings (3-shots, 5-shots respectively) and preset prompts.
-Appendix D Code Repositories
+Appendix F Code Repositories
 We use the Brain-Score repository to evaluate brain alignment for the Pereira2018 and Blank2014 datasets, as well as behavioral alignment for the Futrell2018 dataset.
@@
 Link: www.github.com/awwkl/brain_language_summarization, which builds on www.github.com/mtoneva/brain_language_nlp.
-We use Instruct-Eval repository to evaluate MMLU and BBH scores. Link: www.github.com/declare-lab/instruct-eval.
-We use Stanford Alpaca repository for instruction-tuning. Link: www.github.com/tatsu-lab/stanford_alpaca).
-Appendix E Results for Brain alignment
-Table 5: Brain alignment results for all 25 vanilla and instruction-tuned LLMs. We provide these results for reproducibility purposes.
+We use the Instruct-Eval repository to evaluate MMLU and BBH scores. Link: www.github.com/declare-lab/instruct-eval.
+We use the Stanford Alpaca repository for instruction-tuning. Link: www.github.com/tatsu-lab/stanford_alpaca).
+Appendix G Results for Brain alignment
 Pereira2018 Blank2014 Wehbe2014 Average
@@
 vicuna-33b 0.436 0.156 0.105 0.232
+gpt2-small 0.305 0.123 0.088 0.172
+gpt2-small-alpaca 0.298 0.121 0.081 0.167
+gpt2-medium 0.329 0.081 0.089 0.166
+gpt2-medium-alpaca 0.325 0.080 0.090 0.165
+gpt2-large 0.342 0.077 0.101 0.173
+gpt2-large-alpaca 0.336 0.074 0.101 0.170
+gpt2-xl 0.358 0.140 0.102 0.200
+gpt2-xl-alpaca 0.343 0.139 0.093 0.192
+Table 5: Brain alignment results for all vanilla and instruction-tuned LLMs. We provide these results for reproducibility purposes.
+Pereira2018 Blank2014 Wehbe2014 Average
+Noise ceiling 0.359 0.210 0.104 0.224
 Table 6: Noise ceiling estimates for all 3 neural datasets.
+fMRI measurements inherently include noise, i.e., fluctuations not due to neurons firing, from sources such as environmental factors, the scanner, and movements of human participants.
+Consequently, a “noise ceiling” is often computed by recording multiple samples from the same subjects for the same stimuli, providing an upper-limit estimate of the possible correlation in fMRI activity.
+In our study, we followed prior work for calculating the noise ceiling.
 For Pereira2018 and Blank2014, noise ceiling estimates are computed using the Brain-Score repository, with details provided in Schrimpf et al.
 (2021).
-For Wehbe2014, noise ceiling estimates are also computed using a similar procedure.
-Pereira2018 Blank2014 Wehbe2014 Average
-Noise ceiling 0.359 0.210 0.104 0.224
-Appendix F Results for Next-word prediction, MMLU, BBH
-Table 7: WikiText-2 NWP loss, MMLU Overall Score, and BBH Overall Score for all instruction-tuned LLMs.
-Results for vanilla LLMs are not shown as they are not adapted for the question formats in the MMLU and BBH benchmarks.
-We provide these results for reproducibility purposes.
-WikiText-2 NWP Loss MMLU Overall Score BBH Overall Score
+For Wehbe2014, noise ceiling estimates are computed using a similar procedure.
+Appendix H Results for Next-word prediction, MMLU, BBH
+WikiText-2 NWP Loss MMLU Score BBH Score
 flan-t5-small 0.851 0.294 0.287
@@
 vicuna-33b 3.940 0.590 0.426
+gpt2-small-alpaca 4.193 0.270 0.273
+gpt2-medium-alpaca 3.333 0.260 0.276
+gpt2-large-alpaca 3.614 0.260 0.278
+gpt2-xl-alpaca 3.283 0.270 0.284
+Table 7: WikiText-2 NWP loss, MMLU Overall Score, and BBH Overall Score for all instruction-tuned LLMs.
+Results for vanilla LLMs are not shown as they are not adapted for the question formats in the MMLU and BBH benchmarks.
+Results in gray are close to random performance.
+We provide these results for reproducibility purposes.
 Notes on comparing next-word prediction (NWP) loss across model families.
@@
 (1) Architecture: T5 models have an encoder-decoder architecture while LLaMA models have a decoder-only architecture.
-(2) Training Objectives: The T5 models were trained on supervised and unsupervised tasks, while the LLaMA models were trained only on unsupervised text 3.
+(2) Training Objectives: The T5 models were trained on supervised and unsupervised tasks, while the LLaMA models were trained only on unsupervised text (Section 3).
 (3) Loss computation: The loss functions for both model families are computed differently, making it inappropriate to directly compare their loss values.
@@
 In summary, while NWP loss is a valuable metric for evaluating language models within the same family or architecture, comparing across different model families may not be meaningful.
-Appendix G Results for Correlations of Brain Alignment with LLM properties
-Figure 4: Correlation between brain alignment and various LLM properties: (A) MMLU benchmark global score, (B) BBH benchmark score with only world knowledge tasks, (C) number of parameters of the model, and (D) Next word prediction (NWP) performance.
-Figure 5: Correlation between brain alignment and various LLM properties for all 25 LLMs: (A) number of parameters of the model, and (B) Next word prediction (NWP) performance.
-Appendix H Results for Behavioral alignment
-Table 8: Behavioral alignment results for all 25 vanilla and instruction-tuned LLMs. We provide these results for reproducibility purposes.
+Appendix I Results for Correlations of Brain Alignment with LLM properties
+Figure 6: Correlation between brain alignment and various LLM properties: (A) MMLU benchmark overall score, (B) BBH benchmark score for world knowledge tasks, (C) number of parameters of the model, and (D) Next-word prediction (NWP) loss.
+Figure 7: Correlation between brain alignment and various LLM properties for all LLMs (including vanilla models): (A) number of parameters of the model, and (B) Next-word prediction (NWP) loss.
+Appendix J Additional Brain Alignment results (Linear, RSA, CKA) for Gemma and LLaMA-2 models
+Pretrained models
+Model Linear CKA RSA
+Gemma-2B 0.76 1.94 0.63
+Gemma-7B 0.90 2.25 0.62
+LLaMA-2-7B 0.97 1.81 0.65
+LLaMA-2-13B 1.08 2.51 0.68
+Instruction-tuned models
+Model Linear CKA RSA
+Gemma-2B 0.72 2.44 0.70
+Gemma-7B 0.94 2.55 0.72
+LLaMA-2-7B 1.02 2.00 0.63
+LLaMA-2-13B 1.08 2.77 0.65
+Table 8: Additional Brain Alignment results (Linear, RSA, CKA) for Gemma and LLaMA-2 models.
+“Linear” refers to the linear predictivity metric used in the main paper, normalized by the noise ceiling computed in Appendix G.
+In Section 4, we study the representational similarity between LLMs and human brain activity (brain alignment) using the linear predictivity similarity metric (“Linear”).
+This metric involves training a linear regression model to predict fMRI brain activity based on LLM representations.
+Here, we additionally validate our results using the RSA and CKA similarity metrics (Table 8), confirming that the brain alignment results are consistent across different similarity metrics.
+In Section 4.1, we observed that instruction-tuning generally improves brain alignment for models from the T5 and LLaMA-1 families.
+We further validate this trend with recently released models from the Gemma and LLaMA-2 families (Table 8).
+These results hold across various similarity metrics, including Linear predictivity, RSA, and CKA.
+Appendix K Results for Instruction-tuning LLaMA-7B on Alpaca dataset
+Figure 8: Improvements in brain alignment from instruction-tuning are due to both additional training data, as well as training to understand and follow instructions.
+Instruction model.
+We instruction-tune LLaMA-7B on the Stanford Alpaca dataset (Taori et al., 2023) using the default training process, following the code in www.github.com/tatsu-lab/stanford_alpaca.
+In particular, the model is instruction-tuned using $52$K instruction-following examples generated through methods inspired by Self-Instruct (Wang et al., 2022a).
+This model is labeled “Instruction” in Figure 8.
+No instruction model (Ablation).
+We also train an ablation model with the same process and training data as the default instruction-tuning, but remove the instruction portion from each training sample.
+This ablation model is labeled “No instruction” in Figure 8.
+This ablation experiment disentangles: (1) training data (present in both “Instruction” and “No instruction”), from (2) training LLMs to understand and follow instructions (present only in “Instruction”).
+We use all provided training samples from the Alpaca dataset, thus ensuring that the models are trained on the same amount of data.
+We observe that brain alignment of the “No Instruction” ablation model increases during fine-tuning but stays lower than its “Instruction” counterpart.
+This shows that brain alignment improvements are due to both (1) training data (present in both models) and (2) the process of training LLMs to understand and follow instructions (present only in the “Instruction” model).
+Appendix L Results for Behavioral alignment
 Futrell2018
@@
 vicuna-33b 0.164
-Table 9: Noise ceiling estimates for the Futrell2018 reading-times dataset. Noise ceiling estimates are computed using the Brain-Score repository, with details provided in Schrimpf et al. (2021).
+gpt2-small 0.367
+gpt2-small-alpaca 0.335
+gpt2-medium 0.350
+gpt2-medium-alpaca 0.345
+gpt2-large 0.331
+gpt2-large-alpaca 0.338
+gpt2-xl 0.318
+gpt2-xl-alpaca 0.336
+Table 9: Behavioral alignment results for all vanilla and instruction-tuned LLMs. We provide these results for reproducibility purposes.
 Futrell2018
 Noise ceiling 0.858
-Appendix I Results for Instruction-tuning LLaMA-7B on Alpaca dataset
-Figure 6: Improvements in brain alignment from instruction-tuning are due to both additional training data, as well as training to understand and follow instructions.
-Instruction model
-We instruction-tune LLaMA-7B on the Stanford Alpaca dataset (Taori et al., 2023) using the default training process, following the code in www.github.com/tatsu-lab/stanford_alpaca.
-In particular, the model is instruction-tuned using 5252K instruction-following examples generated through methods inspired by Self-Instruct (Wang et al., 2022a)).
-This model is labeled “Instruction” in Figure 6.
-No instruction model (Ablation)
-We also train an ablation model with the same process and training data as the default instruction-tuning, but remove the instruction portion from each training sample.
-This ablation model is labeled “No instruction” in Figure 6.
-This ablation experiment disentangles: (1) training data (present in both Instruction and No instruction), from (2) training LLMs to understand and follow instructions (present only in Instruction).
-We use all provided training samples from the Alpaca dataset, thus ensuring that the models are trained on the same amount of data.
-We observe that brain alignment of the “No Instruction” ablation model increases during fine-tuning but stays lower than its “Instruction” counterpart.
+Table 10: Noise ceiling estimates for the Futrell2018 reading-times dataset. Noise ceiling estimates are computed using the Brain-Score repository, with details provided in Schrimpf et al. (2021).
+Appendix M Results for Correlations of Behavioral Alignment with LLM properties
+Figure 9: Correlation between behavioral alignment and various LLM properties: (A) MMLU benchmark overall score, (B) number of parameters of the model, (C) Next-word prediction (NWP) loss for T5 models, and (D) NWP loss for LLaMA models.